### v3

Abstract
Diffusion models have recently emerged as a powerful new family of generative models. Remarkably, they are capable of learning image distributions to generate high-quality images that are distinct from the training data, even when trained on a finite dataset. However, worst-case analyses have shown that learning arbitrary distributions via diffusion models may require exponentially many samples in terms of the ambient data dimension, suggesting a gap between theory and practice. In this work, we close this gap by studying how diffusion models learn low-dimensional distributions, motivated by the observation that real-world image datasets often exhibit intrinsic low-dimensional structures. Specifically, we focus on learning a mixture of low-rank Gaussians, where the low-rank covariance captures the underlying low dimensionality of the data and the latent Gaussian structure ensures theoretical tractability. Theoretically, under an appropriate network parameterization, we rigorously show that optimizing the training loss of diffusion models is equivalent to solving the canonical subspace clustering problem over the training samples. Based on this equivalence, we further demonstrate that the minimal number of samples required to learn the distribution scales linearly with the intrinsic dimension under our data and model assumptions. This insight sheds light on why diffusion models can break the curse of dimensionality and exhibit phase transitions in learning distributions. Empirically, we validate our results with various experiments on synthetic and real-world image datasets. Moreover, based on our theoretical findings, we establish a correspondence between the subspace bases and semantic attributes of image data, providing a foundation for controllable image editing. The code is available at https://github.com/huijieZH/Diffusion-Model-Generalizability.
Key words: Diffusion models, mixture of low-rank Gaussians, denoising autoencoder, phase transition
1 Introduction
Generative modeling is a fundamental task in deep learning that aims to learn the underlying data distribution from the training samples to generate new and realistic data. Towards this goal, diffusion models have recently emerged as a powerful class of generative models, achieving remarkable performance across a wide range of domains, including image generation [1], video synthesis [2], and speech and audio generation [3, 4]. In general, diffusion models learn a data distribution from training samples through a process that imitates the nonequilibrium thermodynamic diffusion process [1, 5, 6]. Specifically, diffusion models consist of two stages: (i) a forward diffusion process in which Gaussian noise is incrementally added to training samples at each time step, and (ii) a backward sampling process in which the noise is progressively removed using a neural network trained to approximate the score function of the data distribution, that is, the gradient of the logarithm of the probability density function (pdf) at each time step.
Despite the significant empirical success and recent advances in understanding sampling convergence [7, 8, 9], distribution approximation [10, 11], memorization [12, 13, 14, 15], and generalization [16, 17, 18] of diffusion models, the mechanisms underlying their performance remain poorly understood. This is mainly owing to the black-box nature of neural networks and the inaccessibility of real-world data distributions. In particular, a fundamental question arises: can diffusion models truly learn the underlying data distribution, and if so, how many samples are required to do so? Recent theoretical studies [11, 19] have shown that learning a general probability distribution using diffusion models inevitably suffers from the curse of dimensionality. Specifically, if the underlying density belongs to a generic class of probability distributions, obtaining an $\epsilon$-accurate estimator of the corresponding true score function requires the number of training samples that scales as $O(\epsilon^{-n})$, where $n$ is the ambient data dimension. However, recent empirical studies [16, 18] have demonstrated that diffusion models can successfully learn image data distributions and generate new and sensible samples that are distinct from the training samples, even when trained on a finite dataset. As such, the gap between theory and practice raises a key question:
When and why can diffusion models learn data distributions without suffering from the curse of dimensionality?
1.1 Our Contributions
This work addresses the above question by investigating how diffusion models effectively learn data distributions with low-dimensional intrinsic structures. As opposed to focusing on arbitrary distributions, our study is motivated by the observation that real-world image data often lie on a union of low-dimensional manifolds despite their high ambient dimensions [20, 21]. These structures arise from underlying symmetries, repetitive patterns, and local regularities [22, 23]. To effectively capture this low-dimensional structure of real-world image data while offering analytical tractability, we investigate how diffusion models learn a mixture of low-rank Gaussians (MoLRG; see Definition 1). Notably, our focus on MoLRG is further motivated by promising empirical evidence in Figure 1, which demonstrates that samples generated via the diffusion reverse sampling process—using our theoretically constructed model—closely resemble those produced by the U-Net [24] trained on the same dataset and initialized with the same noise.
Theoretically, we show that diffusion models can learn the MoLRG distribution, provided that the number of training samples scales linearly with the intrinsic dimension of the data. Our result is established by demonstrating an equivalence between the training loss of diffusion models and the canonical subspace clustering problem [25, 26] (see Theorem 1 and Theorem 3) under an appropriate parameterization of the denoising autoencoder. Our theory reveals a phase transition from failure to success in learning the underlying distributions, as the sample size exceeds the intrinsic dimensionality of the subspaces (see Theorem 2 and Theorem 4). Moreover, our theoretical analysis offers valuable practical insights as highlighted below.
  • The phase transition of generalization on image datasets. As shown in Section 4.1, when training diffusion models on real-world image datasets, we observe a similar phase transition from failure to success in generalization, where the model begins to generate new and sensible images distinct from the training data once the number of training samples exceeds a threshold that scales linearly with the intrinsic dimension. Our study of MoLRG offers key insights into understanding this phenomenon in practice.
  • Correspondence between subspace bases and semantic task vectors [^1] A semantic task vector is a direction in the latent or intermediate feature space such that traversing along it causes a controlled and interpretable change in the output.. The subspace basis vectors identified through our theoretical analysis align with semantically meaningful directions—i.e., task vectors—in pretrained diffusion models trained on real-world image datasets. These semantic task vectors enable control over attributes such as gender, hairstyle, and color in the generated images (see Figure 5). This insight has inspired new training-free image editing methods using pretrained diffusion models [27].
1.2 Relationship to Prior Arts
Our study of distribution learning is highly related to recent studies on the generalization of diffusion models. In classical machine learning theory, accurately learning the underlying distribution typically guarantees good generalization [28]. This principle also applies to diffusion models: when they successfully learn the true underlying data distribution, they can generate new samples distinct from the training data, indicating strong generalization capability. On the other hand, recent empirical studies [18, 16] have shown that different diffusion models can reproduce each other’s outputs while generating new samples distinct from the training data, even when trained with different architectures, loss functions, and non-overlapping subsets of the training data. This reproducibility suggests that diffusion models capture the underlying data distribution when they generalize well. Motivated by these discussions, generalization in diffusion models can be viewed as their ability to accurately capture the underlying data distribution.
Extensive recent studies [16, 18] have empirically revealed that diffusion models learn the score function across two distinct regimes–memorization (i.e., learning the empirical distribution) and generalization (i.e., learning the underlying distribution). In the memorization regime, recent studies [18, 29] have shown that diffusion models learn the empirical optimal denoiser when the network is highly over-parameterized. In the generalization regime, extensive works [30, 31] have studied the generalization performance of diffusion models in different settings. Notably, most existing studies on the memorization and generalization of diffusion models are empirical. In contrast, our work provides a rigorous theoretical framework that explains these intriguing experimental observations based on the MoLRG model. Our analysis demonstrates that diffusion models learn the underlying data distribution with minimal sample complexity, which scales linearly with the intrinsic dimension of the data distribution. Our theory reveals a phase transition from failure to success in learning the underlying distribution as the number of training samples increases, shedding light on the phase transition from memorization to generalization.
1.3 Notation
We write matrices in bold capital letters like $\bm{A}$, vectors in bold lower-case letters like $\bm{a}$, and scalars in plain letters like $a$. Given a matrix, we use $\|\bm{A}\|$ to denote its largest singular value (i.e., spectral norm), $\sigma_{i}(\bm{A})$ its $i$-th largest singular value, and $a_{ij}$ its $(i,j)$-th entry, $\mathrm{rank}(\bm{A})$ its rank, $||\bm{A}||_{F}$ its Frobenius norm. Given a vector $\bm{a}$, we use $\|\bm{a}\|$ to denote its Euclidean norm, and $a_{i}$ its $i$-th entry. Let $\mathcal{O}^{n\times d}$ denote the set of $n\times d$ orthonromal matrices.
Organization.
In Section 2, we introduce the preliminaries on diffusion models and some assumptions on the data and model. In Section 3, we present the main results of this paper. In Section 4, we conduct numerical experiments to support our theory and demonstrate its practical implications. Finally, in Section 5, we summarize our work and discuss potential directions for future research.
2 Problem Setup
In this section, we introduce the basics of diffusion models and assumptions regarding the data and models. Here, we consider a training dataset $\{\bm{x}^{(i)}\}_{i=1}^{N}\subseteq\mathbb{R}^{n}$, where each data point is independently and identically distributed (i.i.d.) and sampled according to the underlying data distribution $p_{\rm data}(\bm{x})$, i.e., $\bm{x}^{(i)}\overset{i.i.d.}{\sim}p_{\rm data}(\bm{x})$.
2.1 Preliminaries on Score-Based Diffusion Models
Forward and reverse processes of diffusion models.
In general, diffusion models aim to learn a data distribution and generate new samples through forward and reverse processes indexed by a continuous time variable $t\in[0,1]$. Specifically, the forward process progressively injects noise into the data, which can be described by the following stochastic differential equation (SDE):
$\mathrm{d}\bm{x}_{t}=f(t)\bm{x}_{t}\mathrm{d}t+g(t)\mathrm{d}\bm{w}_{t},$ (1)
where $\bm{x}_{0}\sim p_{\rm data}(\bm{x})$, scalar functions $f(t),g(t):\mathbb{R}\to\mathbb{R}$ denote the drift and diffusion coefficients, respectively, and $\{\bm{w}_{t}\}_{t\in[0,1]}$ is the standard Wiener process. For ease of exposition, let $p_{t}(\bm{x})$ denote the pdf of $\bm{x}_{t}$ and $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ be the transition kernel from $\bm{x}_{0}$ to $\bm{x}_{t}$. [^2] Note that $p_{0}:=p_{\rm data}$. According to (1), one can verify that
$p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}\left(\bm{x}_{t};s_{t}\bm{x}_{0},s^{2}_{t}\sigma_{t}^{2}\bm{I}_{n}\right),$ (2)
where $s_{t}=\exp\left(\int_{0}^{t}f(\xi)\mathrm{d}\xi\right)$ and $\sigma_{t}=\sqrt{\int_{0}^{t}{g^{2}(\xi)}/{s^{2}({\xi})}\mathrm{d}\xi}.$ [^3] With a slight abuse of notation, we denote $s_{t}:=s(t)$ and $\sigma_{t}:=\sigma(t).$ The reverse process gradually removes the noise from $\bm{x}_{1}$ using the following reverse-time ordinary differential equation (ODE):
$\mathrm{d}\bm{x}_{t}=\left(f(t)\bm{x}_{t}-\frac{1}{2}g^{2}(t)\nabla\log p_{t}(\bm{x}_{t})\right)\mathrm{d}t,$ (3)
Note that if $\bm{x}_{1}$ and $\nabla\log p_{t}$ for all $t\in[0,1]$ are known, the reverse process has the same distribution as the forward process at each time $t\geq 0$ [32].
Training loss of diffusion models.
Unfortunately, the score function $\nabla\log p_{t}$ for each $t\in[0,1]$ is usually unknown, because it depends on the underlying data distribution $p_{\rm data}$. To enable data generation via the reverse ODE in (3), diffusion models must estimate the score function $\nabla\log p_{t}$. In addition, Tweedie’s formula [33] establishes an equivalence between the score function $\nabla\log p_{t}(\bm{x}_{t})$ and the posterior mean $\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]$ as follows:
$s_{t}\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]=\bm{x}_{t}+s_{t}^{2}\sigma_{t}^{2}\nabla\log p_{t}(\bm{x}_{t}).$ (4)
This allows us to estimate the posterior mean $\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]$ as an alternative approach for estimating the score function $\nabla\log p_{t}$. Leveraging the strong function approximation capabilities of neural networks [34], recent studies [16, 35] have explored training a time-dependent neural network $\bm{x}_{\bm{\theta}}(\cdot,t):\mathbb{R}^{n}\times[0,1]\to\mathbb{R}^{n}$ with parameters $\bm{\theta}$, referred to as the denoising autoencoder (DAE), to estimate the posterior mean $\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]$. To learn the network parameters $\bm{\theta}$, we minimize the following empirical loss over the training data $\{\bm{x}^{(i)}\}_{i=1}^{N}$:
$\min_{\bm{\theta}}\ell(\bm{\theta})=\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{1}\lambda_{t}\mathbb{E}_{\bm{\epsilon}}$ $\left[\left\|\bm{x}_{\bm{\theta}}(s_{t}\bm{x}^{(i)}+\gamma_{t}\bm{\epsilon},t)-\bm{x}^{(i)}\right\|^{2}\right]\mathrm{d}t,$ (5)
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, $\gamma_{t}:=s_{t}\sigma_{t}$, and $\lambda_{t}:[0,1]\to\mathbb{R}^{+}$ is a weighting function.
2.2 Mixture of Low-Rank Gaussians
Because the distribution of real-world image datasets is generally unknown and difficult to characterize, quantitatively studying the distribution learning behavior of diffusion models on these distributions can be challenging. Alternatively, this work considers learning a mixture of low-rank Gaussians (MoLRG) distribution, which captures the essential low-dimensional structure of real-world image datasets while remaining analytically tractable. Specifically, the MoLRG model is defined as follows.
Definition 1 (Mixture of Low-Rank Gaussians).
We say that a random vector $\bm{x}_{0}\in\mathbb{R}^{n}$ follows a mixture of $K$ low-rank Gaussian distributions with parameters $\{\pi_{k}\}_{k=1}^{K}$, $\{\bm{\mu}_{k}^{\star}\}_{k=1}^{K}\subseteq\mathbb{R}^{n}$, and $\{\bm{\Sigma}^{\star}_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{n\times n}$ if we have
$\bm{x}_{0}\sim\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\bm{\mu}_{k}^{\star},\bm{\Sigma}_{k}^{\star}),$ (6)
where $\pi_{k}\geq 0$ is the mixing proportion of the $k$-th mixture component satisfying $\sum_{k=1}^{K}\pi_{k}=1$, and $\bm{\mu}^{\star}_{k}$ and $\bm{\Sigma}_{k}^{\star}\succeq\bm{0}$ denote the mean and the covariance matrix of the $k$-th component, respectively. In particular, we assume that the covariance matrix $\bm{\Sigma}_{k}^{\star}$ is low-rank with $\mathrm{rank}(\bm{\Sigma}_{k}^{\star})=d_{k}<n$.
Remarks.
Intuitively, data drawn from a MoLRG distribution lie on a union of low-dimensional linear subspaces, where the $k$-th subspace is characterized by the mean $\bm{\mu}_{k}^{\star}$ and the low-rank covariance matrix $\bm{\Sigma}_{k}^{\star}$. Now, we discuss the motivation for studying the model and its connections to other distributions that have been theoretically analyzed.
  • MoLRG  captures the low-dimensional structure of real-world image datasets. Recent studies [20, 21] conducted extensive numerical experiments and demonstrated that image datasets such as MNIST [36], CIFAR-10 [37], and ImageNet [38] approximately reside on a union of low-dimensional manifolds. Locally, each nonlinear manifold can be effectively approximated using its tangent space (i.e., a linear subspace). Consequently, the MoLRG model, which represents the data as a union of linear subspaces, provides a suitable local approximation for real-world image data distributions. This claim is supported by the empirical studies in Section 2.4. In addition, the simplicity of the low-rank Gaussian in the MoLRG model facilitates theoretical analysis, allowing us to derive the closed-form expression for the posterior estimator at each time step, as shown in Lemma 1. Therefore, studying the MoLRG model is a valuable starting point for theoretical studies on the distribution learning capability of diffusion models.
  • Comparison with recent studies on a mixture of isotropic Gaussians. In comparison, many recent studies have investigated how diffusion models learn a mixture of isotropic Gaussians (MoG), that is, $\bm{\Sigma}_{k}^{\star}=\bm{I}_{n}$ in (6); see, e.g., [39, 40]. These studies mainly focus on learning the means of the Gaussian components, provided that the covariance matrices are fixed as identity matrices. In contrast, our work considers a mixture of low-rank Gaussians, where the key challenge lies in learning the low-rank covariance matrices of the components. As such, the MoLRG model, which captures the inherent low-dimensionality of image datasets [22, 23], offers deeper insight into why diffusion models learn the underlying distributions in practice without suffering from the curse of dimensionality.
Figure 1: Comparison of images generated from the Gaussian, MoLRG, and the distribution learned by diffusion models across different datasets. Each row displays images generated from different distributions using the reverse-time ODE sampler, including the Gaussian, MoLRG, and the distribution learned by U-Net. The columns represent images generated from the same initial noise. The results are shown for four datasets: FashionMNIST (top left ), MNIST (top right), CIFAR-10 (bottom left), and FFHQ (bottom right ).
2.3 Network Parameterization Inspired by MoLRG
To analyze the distribution learning behavior of diffusion models, we need to study the training loss in (5) and characterize its optimal solutions. This analysis critically depends on a suitable parameterization of the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$. In practice, $\bm{x}_{\bm{\theta}}(\cdot,t)$ is typically parameterized by a U-Net architecture [24], which consists of deep nonlinear encoder and decoder networks with skip connections. However, the highly nonlinear structure of U-Net poses significant challenges for theoretical analysis.
To simplify the analysis, we consider a network architecture for $\bm{x}_{\bm{\theta}}(\cdot,t)$ that is a combination of multiple one-layer linear encoders and decoders, weighted by a softmax-like function, as follows:
$\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)=\sum_{k=1}^{K}w_{k}(\bm{x}_{t})\left(\bm{\mu}_{k}+\bm{U}_{k}\bm{D}_{k}\bm{U}_{k}^{T}\left(\frac{\bm{x}_{t}}{s_{t}}-\bm{\mu}_{k}\right)\right),$ (7)
where $\bm{\theta}=\{(\pi_{k},\bm{\mu}_{k},\bm{U}_{k},\bm{\Lambda}_{k})\}_{k=1}^{K}$ denotes the network parameters, $\bm{U}_{k}\in\mathcal{O}^{n\times d_{k}}$ has orthonormal columns, and $\bm{\Lambda}_{k}=\mathrm{diag}(\lambda_{k,1},\dots,\lambda_{k,d_{k}})$ is a diagonal matrix. Additionally, $\bm{D}_{k}$ and $w_{k}(\bm{x}_{t})$ are defined as follows:
$\bm{D}_{k}=\mathrm{diag}\left(\frac{s_{t}^{2}\lambda_{k,1}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,1}},\dots,\frac{s_{t}^{2}\lambda_{k,d_{k}}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,d_{k}}}\right),\ w_{k}(\bm{x})=\frac{\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k},s_{t}^{2}\bm{U}_{k}\bm{\Lambda}_{k}\bm{U}_{k}^{T}+\gamma_{t}^{2}\bm{I}_{n}\right)}{\sum_{l=1}^{K}\pi_{l}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{l},s_{t}^{2}\bm{U}_{l}\bm{\Lambda}_{l}\bm{U}_{l}^{T}+\gamma_{t}^{2}\bm{I}_{n}\right)}.$
Our network architecture in (7) can be viewed as a U-Net-based mixture-of-experts architecture [41], where each expert network is a special U-Net consisting of a linear encoder $\bm{U}_{k}^{T}$ and decoder $\bm{U}_{k}$. These experts are then combined through a learnable weighted summation, allowing the model to adaptively assign weights among components. In addition to the resemblance to U-Net, the parameterization in (7) is well motivated from the following perspectives:
  • Inspired by the posterior mean of MoLRG. As the DAE serves as an estimator of the poster mean (i.e., $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\approx\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$), our network parameterization is inspired by the analytical form of the posterior mean $\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ of MoLRG. The detailed derivation can be found in Lemma 1 of Appendix A. Note that the network parameters $\bm{\theta}$ in (7) are learnable instead of being the ground-truth of the means and covariances of the MoLRG. In Section 3, we will investigate how to learn the network parameters in simplified settings.
  • Meaningful image generation through the parameterization. When the network parameters $\bm{\theta}$ are directly estimated from training data, the experimental results in Section 2.4 demonstrate that our parameterization (7) generates images that are coarsely similar to those produced by a standard U-Net. This demonstrates the practical effectiveness of our network parameterization in real-world tasks and further supports its potential for capturing complex image distributions. More details are provided below.
2.4 Experimental Support for Data & Model Assumptions
As illustrated in Figure 1 and Table 1, we empirically validate the MoLRG assumption and the corresponding network parameterization introduced in (7) for approximating real-world data distributions. In our experiments, we use the distribution learned by U-Net as the benchmark. To quantify the similarity between images generated by (7) and those produced by U-Net, we compute the following metric:
$\frac{1}{M}\sum_{i=1}^{M}\left\|\bm{y}_{1}^{(i)}-\bm{y}_{2}^{(i)}\right\|,$ (8)
where $M$ denotes the number of generated samples, and $\bm{y}^{(i)}_{1}$ and $\bm{y}^{(i)}_{2}$ denote the $i$-th samples generated from the distributions learned by U-Net and the proposed parameterization in (7), respectively. Here, both sets of samples are generated using the reverse-time ODE in (3), initialized from the same noise. Our experimental results in Figure 1 on real-world image datasets—including MNIST [42], FashionMNIST [43], CIFAR-10 [37], and FFHQ [44]—demonstrate that (i) the MoLRG distribution captures the intrinsic low-dimensional structure of image data and (ii) our network parameterization generates images that preserve overall layout and semantics, albeit with less accuracy in fine-grained details. Detailed experimental setups are provided in Appendix C.
FashionMNIST MNIST CIFAR-10 FFHQ
Gaussian 72.62 69.12 33.55 36.75
MoLRG 57.56 62.53 31.29 35.78
Table 1: Distance (defined in (8)) between samples generated from the theoretical network parameterization (based on MoLRG or Gaussian) and those by U-Net.
  • Comparison of the generation between our model and U-Net. First, we compare images generated by U-Net trained on the real-world dataset $\{\bm{x}^{(i)}\}_{i=1}^{N}$ with those generated by our parameterized network in (7), which leverages the means and variances estimated from the same dataset. As illustrated in Figure 1, images generated by the two network parameterizations using the same sampling procedure exhibit substantial visual similarity, especially on simpler datasets such as FashionMNIST and MNIST. This observation supports the validity of our data assumption and confirms the effectiveness of our network parameterization in approximating real-world distributions. On more complex datasets such as CIFAR-10 and FFHQ, while MoLRG struggles to capture fine-grained image details, it still preserves the overall structural characteristics of images generated by U-Net. The loss of fine details indicates a limitation of our model assumption that merits further investigation.
  • Comparison of our model with the single full-rank Gaussian parameterization. To provide a more comprehensive evaluation, we compare our model with a network parameterized according to a single full-rank Gaussian model, as explored in prior studies [45, 46]. As illustrated in Figure 1, the single Gaussian parameterization often results in high intra-class variance and blurred images, particularly on simpler datasets such as FashionMNIST and MNIST (second row of the figure). In contrast, our model based on MoLRG significantly improves generation quality (third row) by leveraging multiple classes to mitigate intra-class variance and employing low-rank covariance structures to suppress high-frequency noise. Moreover, as shown in Table 1, despite employing fewer parameters, our model consistently outperforms the single Gaussian model in terms of the distance to the images generated by U-Net.
3 Sample Complexity Analysis for Learning MoLRG
Building upon the setup introduced in Section 2, we theoretically analyze the sample complexity of learning the MoLRG distribution via diffusion models. Specifically, we show that
The training loss of diffusion models in (5) under our parameterization is equivalent to the canonical subspace clustering problem. The minimal data samples required for learning MoLRG via diffusion models scale linearly with the intrinsic data dimension.
For simplicity, we assume that $\bm{\mu}^{\star}_{k}=\bm{0}$ and $\bm{\Lambda}_{k}^{\star}=\bm{I}_{d_{k}}$ for each $k\in[K]$ in MoLRG. Additionally, since real-world images often contain noise due to factors such as sensor limitations, transition error, and environment conditions, we introduce an additive noise term into the MoLRG  model. Consequently, the training samples are generated according to
$\bm{x}^{(i)}=\bm{U}_{k}^{\star}\bm{a}_{i}+\bm{e}_{i}\;\text{with probability}\;\pi_{k},\;\forall i\in[N],$ (9)
where $\bm{a}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\bm{0},\bm{I}_{d_{k}})$ denotes the linear combination coefficients for the orthonormal basis $\bm{U}_{k}^{\star}\in\mathcal{O}^{n\times d_{k}}$ and $\bm{e}_{i}\in\mathbb{R}^{n}$ is noise for each $i\in[N]$. [^4] We should mention that the signal component of this model exactly satisfies Definition 1 because of $\bm{U}_{k}^{\star}\bm{a}_{i}\sim\mathcal{N}(\bm{0},\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T})$. Notably, since the MoLRG distribution is fully characterized by the first and second-order moments of each degenerate Gaussian component, learning this distribution reduces to estimating the bases $\{\bm{U}_{k}^{\star}\}_{k=1}^{K}$. In the following, we demonstrate that this estimation can be achieved by minimizing the DAE training loss (5) with respect to $\bm{\theta}$.
3.1 A Warm-Up Study: Learning a Single Low-rank Gaussian
To provide the intuition, we begin by introducing our result in a simple setting, where the underlying distribution $p_{\rm data}$ is a single low-rank Gaussian, i.e., $K=1$ in (9). Specifically, we have
$\bm{x}^{(i)}=\bm{U}^{\star}\bm{a}_{i}+\bm{e}_{i},$ (10)
where $\bm{U}^{\star}\in\mathcal{O}^{n\times d}$ denotes an orthonormal basis, $\bm{a}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\bm{0},\bm{I}_{d})$ is coefficients for each $i\in[N]$, and $\bm{e}_{i}\in\mathbb{R}^{n}$ is noise for all $i\in[N]$. According to (7), the optimal parameterization of the DAE in this case reduces to
$\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)=\frac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\bm{U}\bm{U}^{T}\bm{x}_{t},$ (11)
where $\bm{\theta}=\bm{U}\in\mathcal{O}^{n\times d}$. Equipped with the above setup, we show the following result.
Theorem 1.
Suppose that the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$ in Problem (5) is parameterized into (11) for each $t\in[0,1]$. Then, Problem (5) is equivalent to the following principal component analysis (PCA) problem:
$\max_{\bm{U}\in\mathbb{R}^{n\times d}}\sum_{i=1}^{N}\|\bm{U}^{T}\bm{x}^{(i)}\|^{2}\qquad\mathrm{s.t.}\quad\bm{U}^{T}\bm{U}=\bm{I}_{d}.$ (12)
We defer the proof to Appendix B.1. In this case, Theorem 1 shows that training diffusion models with the network parameterization (11) is equivalent to performing PCA on the training samples. Note that PCA is a classical and well-studied method for learning low-dimensional subspaces, whose optimal solution can be computed via singular value decomposition (SVD). This closed-form solution allows us to leverage existing results, such as Wedin’s Theorem [47], to facilitate our analysis. Leveraging this equivalence and using the concentration inequalities, we can precisely determine the minimal number of samples required to learn a single low-rank Gaussian as follows.
Theorem 2.
Consider the setting of Theorem 1. Suppose that the training samples $\{\bm{x}^{(i)}\}_{i=1}^{N}$ are generated according to (10). Let $\hat{\bm{U}}$ denote an optimal solution of Problem (5). The following statements hold:
  • i)
    If $N\geq d$, it holds with probability at least $1-1/2^{N-d+1}-\exp\left(-c_{2}N\right)$ that any optimal solution $\hat{\bm{U}}$ satisfies
    $\left\|\hat{\bm{U}}\hat{\bm{U}}^{T}-\bm{U}^{\star}\bm{U}^{\star T}\right\|_{F}\leq\frac{c_{1}\sqrt{\sum_{i=1}^{N}\|\bm{e}_{i}\|^{2}}}{\sqrt{N}-\sqrt{d-1}},$ (13)
    where $c_{1},c_{2}>0$ are constants.
  • ii)
    If $N<d$, there exists an optimal solution $\hat{\bm{U}}\in\mathcal{O}^{n\times d}$ such that with probability at least $1-1/2^{d-N+1}-\exp\left(-c_{2}^{\prime}d\right)$,
    $\left\|\hat{\bm{U}}\hat{\bm{U}}^{T}-\bm{U}^{\star}\bm{U}^{\star T}\right\|_{F}\geq\alpha-\frac{c_{1}^{\prime}\sqrt{\sum_{i=1}^{N}\|\bm{e}_{i}\|^{2}}}{\sqrt{d}-\sqrt{N-1}},$ (14)
    where $\alpha:=\sqrt{2\min\{d-N,n-d\}}$ and $c_{1}^{\prime},c_{2}^{\prime}>0$ are constants .
(a) PCA
(b) Diffusion Model
Figure 2: Phase transition of learning the MoLRG distribution with $K=1$. The $x$-axis is the number of training samples and $y$-axis is the dimension of subspaces. Darker pixels represent a lower empirical probability of success. We apply SVD and train diffusion models to solve Problems (12) and (5), visualizing the results in (a) and (b), respectively.
We defer the proof to Appendix B.2. Here, we make some remarks on the implications of our result.
  • Phase transition in learning the underlying distribution. Building on the equivalence in Theorem 1 and the data model in (10), Theorem 2 clearly demonstrates a phase transition from failure to success of diffusion models for learning the underlying distribution as the number of training samples increases. More precisely, when the number of training samples is larger than the intrinsic dimension of the subspace, i.e., $N\geq d$, any optimal solution $\hat{\bm{U}}$ recovers the basis of the underlying subspace with an approximation error depending on the noise level. Conversely, when $N<d$, optimizing the training loss fails to learn the underlying distribution. This phase transition is further corroborated by our experiments in Figures 2. Finally, since a Gaussian distribution is fully characterized by its first two moments, our results rigorously show that diffusion models can recover the underlying distribution when $N\geq d$, with the covariance approximation error bounded by the noise level.
  • The connections between PCA and semantic task vectors. The correspondence between principal components and semantic meaning has been well studied in machine learning literature. For example, early work [48] demonstrated that PCA can uncover meaningful axes of variation in natural image datasets, such as facial expressions, lighting, or pose, suggesting a connection between directions of maximal variance and human-perceived semantics. This finding laid the groundwork for interpreting principal components as capturing dominant semantic directions in diffusion models. To further support this perspective, we present experimental evidence in Section 4.2.
3.2 Learning a Mixture of Low-Rank Gaussians
In this subsection, we extend the above study to the MoLRG distribution with $K>1$. For ease of analysis, we assume that the basis matrices of subspaces satisfy $\bm{U}_{k}^{\star T}\bm{U}_{l}^{\star}=\bm{0}$ for each $k\neq l$, $d_{1}=\dots=d_{K}=d$, and the mixing weights satisfy $\pi_{1}=\dots=\pi_{K}=1/K$. Moreover, we consider a hard-max counterpart of (7) for the DAE parameterization as follows:
$\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)=\dfrac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\bm{U}\bm{\Lambda}\bm{U}^{T}\bm{x}_{t}=\dfrac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\sum_{k=1}^{K}\hat{w}_{k}(\bm{\theta};\bm{x}_{0})\bm{U}_{k}\bm{U}_{k}^{T}\bm{x}_{t},$ (15)
where $\bm{\theta}=\{\bm{U}_{k}\}_{k=1}^{K}$, $\bm{U}=[\bm{U}_{1},\dots,\bm{U}_{K}]\in\mathcal{O}^{n\times\sum_{k=1}^{K}d_{k}}$, $\bm{\Lambda}=\mathrm{diag}\left(\hat{w}_{1}(\bm{\theta};\bm{x}_{0}),\dots,\hat{w}_{K}(\bm{\theta};\bm{x}_{0})\right)$, and $\{\hat{w}_{k}(\bm{\theta};\bm{x}_{0})\}_{k=1}^{K}$ are set as the following hard-max weights:
$\hat{w}_{k}(\bm{\theta};\bm{x}_{0})=\begin{cases}1,\ &\text{if}\ k=k_{0},\\ 0,\ &\text{otherwise},\end{cases}$ (16)
where $k_{0}\in[K]$ is an index satisfying $\|\bm{U}_{k_{0}}^{T}\bm{x}_{0}\|\geq\|\bm{U}_{l}^{T}\bm{x}_{0}\|$ for all $l\neq k_{0}\in[K]$. We refer the reader to Appendix B.3 for a discussion on using hard-max weights to approximate the soft-max weights in (7). Now, we are ready to show the following theorem.
(a) Subspace Clustering
(b) Diffusion Model
Figure 3: Phase transition of learning the MoLRG distribution with $K=2$. The $x$-axis is the number of training samples and $y$-axis is the dimension of subspaces. Darker pixels represent a lower empirical probability of success. We apply a subspace clustering method and train diffusion models for solving Problems (17) and (5), visualizing the results in (a) and (b), respectively. Additional experiments for the case when $K=3$ are presented in Figure 6.
Theorem 3.
Suppose that the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$ in Problem (5) is parameterized into (15) for each $t\in[0,1]$, where $\hat{w}_{k}(\bm{\theta},\bm{x}_{0})$ is defined in (16) for each $k\in[K]$. Then, Problem (5) is equivalent to the following subspace clustering problem:
$\max_{\bm{\theta}}\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}(\bm{\theta})}\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}\qquad\mathrm{s.t.}\quad\left[\bm{U}_{1},\dots,\bm{U}_{K}\right]\in\mathcal{O}^{n\times dK},$ (17)
where $C_{k}(\bm{\theta}):=\left\{i\in[N]:\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|\geq\|\bm{U}_{l}^{T}\bm{x}^{(i)}\|,\ \forall l\neq k\right\}$ for each $k\in[K]$.
Here, the constraint set $C_{k}(\bm{\theta})$ ensures that each data point is assigned to the subspace onto which the norm of its projection is largest. Problem (17) seeks to maximize the sum of squared norms of the projections of data points onto their respective assigned subspaces. We defer the proof of this theorem to Appendix B.4. With the network parameterization in (15), Theorem 3 shows that optimizing the training loss of diffusion models is equivalent to solving the subspace clustering problem [25, 26]. Notably, subspace clustering is a fundamental problem in unsupervised learning, which aims to identify and group data points that lie in a union of low-dimensional subspaces in a high-dimensional space [25, 49]. It has been extensively studied both algorithmically and theoretically in the literature; see, e.g., [50, 26]. By showing an equivalence between training diffusion models and subspace clustering in Theorem 3, it enables us to characterize the minimum number of samples required for learning the underlying MoLRG distribution similar to Theorem 2.
Theorem 4.
Consider the setting of Theorem 3. Suppose that the training samples $\{\bm{x}^{(i)}\}_{i=1}^{N}$ are generated according to (9), where $\bm{U}_{k}^{\star T}\bm{U}_{l}^{\star}=\bm{0}$ for each $k\neq l$, $d_{1}=\dots=d_{K}=d$, and $\pi_{1}=\dots=\pi_{K}=1/K$. Suppose in addition that $d\gtrsim\log N$ and $\|\bm{e}_{i}\|\lesssim\sqrt{d/N}$ for all $i\in[N]$. Let $\{\hat{\bm{U}}_{k}\}_{k=1}^{K}$ denote an optimal solution of Problem (5) and $N_{k}$ denote the number of samples from the $k$-th Gaussian component. Then, the following statements hold:
  • (i)
    If $N_{k}\geq d$ for each $k\in[K]$, there exists a permutation $\Pi:[K]\to[K]$ such that with probability at least $1-2K^{2}N^{-1}-\sum_{k=1}^{K}\left(1/2^{N_{k}-d+1}+\exp\left(-c_{2}N_{k}\right)\right)$ for each $k\in[K]$,
    $\left\|\hat{\bm{U}}_{\Pi(k)}\hat{\bm{U}}_{\Pi(k)}^{T}-\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T}\right\|_{F}\leq\frac{c_{1}\sqrt{\sum_{i=1}^{N}\|\bm{e}_{i}\|^{2}}}{\sqrt{N}_{k}-\sqrt{d-1}},$ (18)
    where $c_{1},c_{2}>0$ are constants.
  • (ii)
    If $N_{k}<d$ for some $k\in[K]$, there exists a permutation $\Pi:[K]\to[K]$ and $k\in[K]$ such that with probability at least $1-2K^{2}N^{-1}-\sum_{k=1}^{K}\left(1/2^{d-N_{k}+1}+\exp\left(-c_{2}^{\prime}N_{k}\right)\right)$,
    $\left\|\hat{\bm{U}}_{\Pi(k)}\hat{\bm{U}}_{\Pi(k)}^{T}-\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T}\right\|_{F}\geq\beta-\frac{c_{1}^{\prime}\sqrt{\sum_{i=1}^{N}\|\bm{e}_{i}\|^{2}}}{\sqrt{d}-\sqrt{N_{k}-1}},$ (19)
    where $\beta:=\sqrt{2\min\{d-N_{k},n-d\}}$ and $c_{1}^{\prime},c_{2}^{\prime}>0$ are constants.
We defer the proof of this theorem to Appendix B.5. This result generalizes the finding in Theorem 2. Specifically, when the assignments of data points are known, the subspace clustering problem reduces to $K$ independent PCA problems. Moreover, we not only theoretically establish a phase transition in learning the underlying subspaces but also empirically validate this in Figure 3. Now, we discuss the implications of our results below.
  • Understanding diffusion models via subspace clustering. To the best of our knowledge, our work is the first to establish the equivalence between training diffusion models and subspace clustering. This equivalence, together with the MoLRG model, allows us to show that the minimal number of samples for diffusion models to learn the underlying distribution scales linearly with the intrinsic dimension. This finding stands in sharp contrast to existing results [11, 19] in the literature, which show that diffusion models suffer from the curse of dimensionality when learning distributions. Our results offer a more optimistic and practically relevant perspective by showing that diffusion models can learn data distributions with underlying low-dimensional structures without suffering from the curse of dimensionality.
  • Connections to the phase transition from memorization to generalization. In Section 1.2, we have introduced the connection between generalization and distribution learning and the definitions of memorization and generalization in diffusion models. Now, we clarify the difference between the phase transition in distribution learning described in Theorems 2 & 4 and the transition from memorization to generalization [18, 16]. Our phase transition refers to the shift from failure to success of learning the underlying distribution as the number of training samples increases. In contrast, the latter concerns the shift from memorizing data to generalization as the number of training samples increases. Nevertheless, our theory still sheds light on the minimal number of samples required for diffusion models to enter the generalized regime.
  • Future directions based on our theory. There are several promising directions for future research based on our theoretical framework. First, our current analysis assumes that the data lies on a union of mutually orthogonal subspaces. This facilitates theoretical tractability but does not fully capture the complexity of real-world data, which often resides on overlapping or even nonlinear manifolds. Extending our framework to accommodate such more realistic structures would broaden its applicability. Second, our analysis focuses on a simplified network parameterization for learning MoLRG. In contrast, practical diffusion models typically rely on complex and over-parameterized architectures, such as U-Net and Transformers. A compelling direction for future research is to study the generalization behavior of diffusion models under these over-parameterized and practical network architectures.
3.3 Empirical Validation of Theoretical Findings
We conclude this section by providing phase transition experiments for $K=1,2,3$, shown in Figure 2, Figure 3, and Figure 6. Our experimental results show that training diffusion models consistently exhibits a phase transition from failure to success in learning the MoLRG distribution (or the subspaces) as the number of training samples increases, supporting our theoretical findings in Theorems 2 and 4.
Specifically, our experimental setup is as follows. For each plot, we fix the ambient dimension of the data to be $n=48$ and vary the subspace dimension $d$ from $2$ to $8$ in increments of $1$. Similarly, we vary the number of training samples $N$ from $2$ to $15$ in steps of $1$. For each pair of $(n,d)$, we generate all training samples according to the MoLRG distribution in (9) with $\bm{e}_{i}=0$ for different $K=1,2,3$, independently repeating the experiment for $20$ times to empirically estimate the success probability of subspace recovery.
For the case $K=1$, we apply SVD to solve the PCA problem in (12). To solve the subspace clustering problem in (17) when $K>1$, we apply the K-subspace method with spectral initialization as described in [26]. To train the DAE with the theoretical parameterization (11) or (15), we optimize the training loss (5) via stochastic gradient descent (see Algorithm 1 for more details).
4 Practical Implications of Our Theoretical Results
Building on the results in Section 3, we study the practical value of our theoretical investigation by showing that: (i) our study of low-dimensional distribution learning offers key insights into the generalization behavior of real-world diffusion models (see Section 4.1), and (ii) the basis vectors of the identified low-dimensional subspaces correspond to different semantic task vectors in practice, enabling controlled editing of specific attributes in content generation (see Section 4.2).
Figure 4: Phase transition of generalization using U-Net. Diffusion models with a U-Net architecture are trained on synthetic data sampled from the MoLRG distribution (left column; $K=2$, $n=48$, varying intrinsic dimensions) and on real image datasets: CIFAR-10, CelebA, FFHQ, and AFHQ (right column). The GL score is plotted against the ratio of training samples to the intrinsic dimension (top row) and to the square of the intrinsic dimension (bottom row). A black dashed line fits the data across different intrinsic dimensions (datasets) for each figure. A GL score above 0.95 (within the dark grey region) indicates good generalization, while a score below 0.95 ( within the light grey region) indicates poor generalization.
4.1 Phase Transition of Generalization in Real-World Diffusion Models
In this subsection, we conduct experiments on both synthetic MoLRG data and real image datasets to train U-Net-based diffusion models. Consistent with the predictions of Theorem 2 and Theorem 4, we observe a similar phase transition in generalization from failure to success, on both synthetic and real image datasets. As discussed in Section 1.2, achieving good generalization is closely tied to accurately learning the underlying distribution in diffusion models. More specifically, we empirically show that the minimum number of training samples, denoted by $N_{\mathrm{min}}$, required for generalization scales linearly with the intrinsic dimension, denoted by $\mathrm{ID}$, on both synthetic and real datasets,
$N_{\mathrm{min}}=c\cdot\mathrm{ID},$ (20)
where $c>0$ is a constant. Therefore, our theoretical framework not only explains distribution learning in the MoLRG model but also offers valuable insights into the generalization of real-world diffusion models.
Measuring generalization in diffusion models.
As discussed in Section 1.2, recent studies [18, 16] have shown that diffusion models trained under different settings can reproduce each other’s outputs. This reproducibility provides strong evidence of generalization [16]. Furthermore, [18] empirically demonstrates that this phenomenon co-emerges with the models’ ability to generate novel samples distinct from their training data. Taken together, these findings suggest that the ability of diffusion models to generate new samples can serve as an indicator of good generalization. Specifically, let $\{\bm{y}^{(j)}\}_{j=1}^{M}$ denote $M$ samples generated by a diffusion model trained on the dataset $\{\bm{x}^{(i)}\}_{i=1}^{N}$. Then, we adopt a variant of the generalization (GL) score proposed in [18], defined as follows:
$\text{GL}:=\frac{1}{M}\sum_{j=1}^{M}\mathbb{I}\left(\min_{i\in[N]}\|\bm{\Psi}\left(\bm{x}^{(i)}\right)-\bm{\Psi}\left(\bm{y}^{(j)}\right)\|\geq\delta\right).$ (21)
Here, $\delta$ is a pre-defined threshold, $\bm{\Psi}(\bm{x})$ denotes a descriptor function applied to $\bm{x}$, and $\mathbb{I}(\cdot)$ is the indicator function, where $\mathbb{I}(x\geq\delta)=1$ if $x\geq\delta$ and 00 otherwise. For the MoLRG distribution, we set $\bm{\Psi}(\bm{x})$ as the identity function, and $\delta$ is defined in (46) in Appendix E.1. For real-world datasets, we set $\bm{\Psi}(\bm{x})$ as the self-supervised copy detection descriptor introduced in [51], a neural feature extractor tailored for copy detection tasks, and set $\delta=0.8$ according to [51, 52]. Additional details are provided in Appendix E.
Intuitively, the GL score measures the dissimilarity between the generated samples $\{\bm{y}^{(j)}\}_{j=1}^{M}$ and the training samples $\{\bm{x}^{(i)}\}_{i=1}^{N}$ in the feature space. A higher GL score indicates that the generated samples are less similar to the training data, reflecting better generalization. In this work, we consider a diffusion model to generalize well when $\text{GL}>0.95$, i.e., at least 95% of the generated samples are distinct from the training set.
Experiments on synthetic data.
In this subsection, we demonstrate a phase transition in generalization when training U-Net on synthetic data generated from the MoLRG distribution. We use the MoLRG model defined in (9) and set the data dimension $n=48$, the number of components $K=2$, the noise level $\bm{e}_{i}=\bm{0}$, and mixing proportion $\pi_{k}=1/2$. Then, we set each cluster to contain an equal number of samples and the total number is $N$. The intrinsic dimension of each subspace is set to $d$, identical across clusters. Since the subspace bases are orthogonal, the total intrinsic dimension of the distribution $\mathrm{ID}=Kd$. We optimize the training loss in (5) using a DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$, parameterized by U-Net. The same U-Net architecture is used across all experiments. Detailed experimental settings are provided in Appendix E.1.
In Figure 4 (top-left), we plot the GL scores against the ratio $\log_{2}\left(N/\mathrm{ID}\right)$ by varying both $N$ and $\mathrm{ID}$. Here, different scatter colors correspond to different choices of $\mathrm{ID}=8,10,12$. It is observed that the GL scores plotted against $\log_{2}\left(N/\mathrm{ID}\right)$ consistently exhibit a sigmoid-shaped curve across different $\mathrm{ID}$ values. This suggests that, for a fixed model architecture, generalization ability depends primarily on the ratio $N/\mathrm{ID}$ rather than on the values of $N$ or $\mathrm{ID}$ individually. Specifically, we fit all the points using a sigmoid function (the back dashed curve shown in Figure 4 top-left), denoted by $f_{\texttt{MoLRG}}\left(N/\mathrm{ID}\right)$, with details provided in Appendix E.1. For comparison, we plot the GL scores against $\log_{2}\left(N/\mathrm{ID}^{2}\right)$ in Figure 4 (bottom-left) and fit them with a sigmoid function. It is observed that the data point deviates more from the fitted curve compared to the top plot. The stronger alignment between the data points and the fitted curve in the top plot confirms that $N/\mathrm{ID}$ is a better indicator for GL score.
Recall that $\text{GL}>0.95$ indicates successful generalization. To identify $N_{\mathrm{min}}$, we solve $\text{GL}({N_{\mathrm{min}}}/\mathrm{ID})\approx f_{\texttt{MoLRG}}\left(N_{\mathrm{min}}/\mathrm{ID}\right)=0.95$, which implies $c=f_{\texttt{MoLRG}}^{-1}(0.95)$ in (20). [^5] It is worth noting that the linear relationship between $N_{\mathrm{min}}$ and $\mathrm{ID}$ holds as long as the data points (plotting GL score against $N/\mathrm{ID}$) can be well-fitted by a function. Changing the threshold for successful generalization affects only the slope $c$ of the linear relationship. This demonstrates that achieving successful generalization requires $N_{\mathrm{min}}$ to scale linearly with $\mathrm{ID}$, thereby corroborating our theoretical findings in Theorem 4.
Experiments on real image datasets.
Next, our results in Figure 4 (top-right) reveal a similar phase transition in generalization across several real-world image datasets, including CIFAR-10, CelebA [53], FFHQ, and AFHQ [54]. Following a similar experimental setup for MoLRG, we use the same U-Net architecture for different datasets with extra experimental details provided in Appendix E.2. Similar to the study of MoLRG, we respectively plot the GL score against $\log_{2}(N/\mathrm{ID})$ and $\log_{2}(N/\mathrm{ID^{2}})$ by varying the number of training samples $N$. However, because the intrinsic dimension of image datasets here is not known, we estimate it using the method described in Appendix E.3, with the resulting estimates summarized in Table 2.
CIFAR-10 CelebA FFHQ AFHQ
$\mathrm{ID}$ 10.8 11.5 15.8 16.7
Table 2: Intrinsic dimensions $\mathrm{ID}$ for different real world datasets.
Our results across different real image distributions also suggest that the GL score is primarily determined by the ratio $N/\mathrm{ID}$. As shown in Figure 4 (top-right), the plot of the GL score against $\log_{2}\left(N/\mathrm{ID}\right)$ yields nearly identical sigmoid-shaped curves across different datasets. This behavior is consistent with our observations for the MoLRG distribution. In contrast, Figure 4 (bottom-right) shows the GL scores plotted against $N/\mathrm{ID^{2}}$, where data points cannot be captured by a single function across datasets. This comparison further supports that $N/\mathrm{ID}$ is a more appropriate indicator for GL score.
Accordingly, we fit all the points of GL scores using the same function $f_{\texttt{real}}\left(N/\mathrm{ID}\right)$ (the black dashed curve shown in Figure 4 top-right), with more details provided in Appendix E.2. In this case, training diffusion models on real image datasets requires at least $N_{\mathrm{min}}=f^{-1}_{\texttt{real}}\left(0.95\right)\mathrm{ID}$ number of training data to achieve good generalization. This aligns with the linear relationship in (20) with constant $c=f^{-1}_{\texttt{real}}\left(0.95\right)$.
Figure 5: Correspondence between the singular vectors of the Jacobian of the DAE and semantic image attributes. We use a pre-trained DDPM with U-Net on the MetFaces dataset [55]. We edit the original image $\bm{x}_{0}$ by changing $\bm{x}_{t}$ into $\bm{x}_{t}+\alpha\bm{q}_{i}$, where $\bm{q}_{i}$ is a singular vector of the Jacobian of the DAE $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$.
4.2 Correspondence between Basis of Low-Dimensional Subspaces and Semantic Attributes
In this subsection, we demonstrate that our theoretical insight provides valuable guidance for improved controllability in image generation. We begin by outlining a method for identifying low-rank subspaces within diffusion models trained on real-world image datasets. Next, we demonstrate how to verify that the orthogonal basis vectors of the identified subspaces are semantic task vectors. As illustrated in Figure 5, these vectors can be leveraged to steer diffusion models for editing image attributes such as gender, hairstyle, and color. While previous works [56, 57] have explored similar methods for image editing, our study offers a new perspective to understand the method through the lens of a low-dimensional subspace. Building on this study, our concurrent work [27] proposes a training-free method that enables controllable image editing.
Identifying Low-rank Subspaces in Diffusion Models.
Although the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$ of real-world diffusion models cannot be exactly written as the theoretical parameterization introduced in (15), it is still possible to identify a low-rank subspace locally. Recent studies [46, 27] have shown that $\bm{x}_{\bm{\theta}}(\cdot,t)$ can be well approximated using a first-order Taylor expansion:
$\bm{x}_{\bm{\theta}}(\bm{x}_{t}+\bm{\delta}_{t},t)\approx\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)+\bm{J}_{t}\bm{\delta}_{t},$ (22)
where $\bm{\delta}_{t}$ is the steering direction and $\bm{J}_{t}=\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ denotes the Jacobian of the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$ at $\bm{x}_{t}$. As we empirically verify in Appendix E.3, $\bm{J}_{t}$ is often a low-rank matrix at certain timesteps $t$, indicating that its range spans a low-dimensional subspace around $\bm{x}_{t}$. To identify an orthonormal basis for this subspace, we apply an SVD to $\bm{J}_{t}$ and obtain $\bm{J}_{t}=\bm{P}\bm{\Sigma}\bm{Q}^{T}$, where $r:=\mathrm{rank}(\bm{J}_{t})$, $\bm{P}=[\bm{p}_{1},\cdots,\bm{p}_{r}]\in\mathcal{O}^{n\times r}$, $\bm{Q}=[\bm{q}_{1},\cdots,\bm{q}_{r}]\in\mathcal{O}^{n\times r}$, and $\bm{\Sigma}=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})$ with $\sigma_{1}\geq\dots\geq\sigma_{r}\geq 0$. Each $\bm{p}_{i}$ serves as an orthogonal basis vector for the subspace.
As such, if we choose $\bm{\delta}_{t}=\alpha\bm{q}_{i}$ to be one of the singular vector $i\in[r]$, then we obtain
$\bm{x}_{\bm{\theta}}(\bm{x}_{t}+\alpha\bm{q}_{i},t)$ $\approx\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)+\sum_{j=1}^{r}\sigma_{j}\bm{p}_{j}\langle\bm{q}_{j},\alpha\bm{q}_{i}\rangle$
$=\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)+\alpha\sigma_{i}\bm{p}_{i}.$
Given that $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\approx\mathbb{E}[\bm{x}_{0}\mid\bm{x}_{t}]$ serves as an estimate of the clean image, this implies that applying a perturbation $\bm{\delta}=\alpha\bm{q}_{i}$ modifies the original image $\bm{x}_{0}$ along the direction of the corresponding orthogonal basis vector $\bm{p}_{i}$, with an effective strength of $\alpha\sigma_{i}$. In the following, we empirically show that each $\bm{p}_{i}$ is often associated with semantic attributes.
Experimental implementation and results.
Now, we use a pre-trained diffusion denoising probabilistic model (DDPM) [1] on the MetFaces dataset [55]. We randomly select an image $\bm{x}_{0}$ from the dataset and use the reverse process of the diffusion denoising implicit model (DDIM) [58] to generate $\bm{x}_{t}$ at $t=0.7$. We choose the steering direction as the leading right singular vectors $\bm{q}_{1},\ldots\bm{q}_{5}$ and use $\tilde{\bm{x}}_{t}=\bm{x}_{t}+\alpha\bm{q}_{i}$ to generate new images with editing strength $\alpha\in[-6,6]$. Figure 5 shows that these singular vectors enable different semantic edits in terms of gender, hairstyle, and color of the image. For comparison, steering the image along a direction $\bm{s}$ drawn uniformly at random from the unit sphere results in almost no perceptible change in the edited images. This implies that the low-dimensional subspace spanned by $\bm{P}$ is non-trivial, with its leading basis vectors corresponding to semantic task vectors. Experimental results for more images and ablation studies for $t=0.1$ and $0.9$ are shown in Figure 8(b).
5 Conclusion & Future Directions
In this work, we studied the training loss of diffusion models to investigate when and why diffusion models can learn the underlying distribution without suffering from the curse of dimensionality. By assuming the data following MoLRG motivated by extensive empirical observations, we showed that minimizing the training loss of diffusion models is equivalent to solving the subspace clustering problem under proper network parameterization. Based on this equivalence, we further showed that the optimal solutions to the training loss can recover the underlying subspaces when the minimal number of samples scales linearly with the intrinsic dimensionality of the data distribution. Moreover, we established the correspondence between the subspaces and the semantic representations of image data.
Our work opens several new directions for advancing the understanding of the underlying mechanisms of diffusion models. First, as noted in the remarks of Theorem 4, while our work explains the generalization ability of diffusion models, it does not address the phenomenon of memorization or the phase transition from memorization to generalization. Future work is to extend the current analysis to consider over-parametrized models and explore how these models contribute to memorization and generalization. Second, our work mainly focuses on leveraging low-dimensional structures to understand the training process of diffusion models. However, the sampling process is also critical in diffusion models, as it influences the efficiency and quality of generated samples. A key direction for future research is to explore how low-dimensional structures, such as the MoLRG framework, can be utilized to improve the sampling efficiency of diffusion models.
Acknowledgements
P.W., H.Z., Z.Z., S.C., and Q.Q. acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, NSF IIS 2402950, a gift grant from KLA, and the MICDE Catalyst Grant. Y.M. acknowledges support from the joint Simons Foundation-NSF DMS grant #2031899, NSF IIS 2402951, and the ONR grant N00014-22-1-2102. The authors also would like to thank Laura Balzano (U. Michigan), Jeff Fessler (U. Michigan), Huikang Liu (SJTU), Dogyoon Song (UC Davis), Liyue Shen (U. Michigan), Rene Vidal (UPenn), and Zhihui Zhu (OSU) for stimulating discussions.
Appendix A Proofs in Section 2
When the data $\bm{x}_{0}$ is drawn from the MoLRG distribution defined in Definition 1, the simplicity of Gaussian components allows us to derive a closed-form expression for the ground-truth posterior mean $\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]$ for all $t\in[0,1]$ as follows. To proceed, for each $\bm{\Sigma}_{k}^{\star}$, we have its eigen-decomposition as follows:
$\bm{\Sigma}_{k}^{\star}=\bm{U}_{k}^{\star}\bm{\Lambda}_{k}^{\star}\bm{U}_{k}^{\star T},$ (24)
where $\bm{\Lambda}_{k}^{\star}=\mathrm{diag}\left(\lambda_{k,1}^{\star},\dots,\lambda_{k,d_{k}}^{\star}\right)$ is a diagonal matrix with $\lambda_{k,1}^{\star}\geq\dots\geq\lambda_{k,d_{k}}^{\star}>0$ being its positive eigenvalues and $\bm{U}_{k}^{\star}\in\mathcal{O}^{n\times d_{k}}$ is an orthonromal matrix whose columns are the corresponding eigenvectors. We first compute the ground-truth score function as follows:
Proposition 1.
Suppose that the underlying data distribution $p_{\rm data}$ is a mixture of low-rank Gaussian distributions in Definition 1. In the forward process of diffusion models, the pdf of $\bm{x}_{t}$ for each $t>0$ is
$p_{t}(\bm{x})=\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right),$ (25)
where $\gamma_{t}:=s_{t}\sigma_{t}$. Moreover, the score function of $p_{t}(\bm{x})$ is
$\nabla\log p_{t}(\bm{x})=\frac{1}{\gamma_{t}^{2}}\frac{\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)\left(\bm{I}_{n}-\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\right)\left(s_{t}\bm{\mu}_{k}^{\star}-\bm{x}\right)}{\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)},$ (26)
where $\bm{D}_{k}^{\star}=\mathrm{diag}\left(\frac{s_{t}^{2}\lambda_{k,1}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,1}^{\star}},\dots,\frac{s_{t}^{2}\lambda_{k,d_{k}}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,d_{k}}^{\star}}\right)$.
Proof.
Let $Y\in\{1,\dots,K\}$ be a discrete random variable that denotes the value of components of the mixture model. Note that $\gamma_{t}=s_{t}\sigma_{t}$. It follows from Definition 1 that $\mathbb{P}(Y=k)=\pi_{k}$ for each $k\in[K]$. Conditioned on $Y=k$, we have $\bm{x}_{0}\sim\mathcal{N}(\bm{\mu}_{k}^{\star},\bm{\Sigma}_{k}^{\star})$. This, together with (2), implies $\bm{x}_{t}\sim\mathcal{N}\left(s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)$. Therefore, we have
$p_{t}(\bm{x})=\sum_{k=1}^{K}p_{t}(\bm{x}|Y=k)\mathbb{P}\left(Y=k\right)=\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right).$
Next, we directly compute
$\nabla\log p_{t}(\bm{x})$ $=\frac{\nabla p_{t}(\bm{x})}{p_{t}(\bm{x})}=\frac{\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)\left(s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)^{-1}\left(s_{t}\bm{\mu}_{k}^{\star}-\bm{x}\right)}{\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)}.$
Using (24) and the matrix inversion lemma, we compute
$\left(s_{t}^{2}\bm{\Sigma}^{\star}_{k}+\gamma_{t}^{2}\bm{I}_{n}\right)^{-1}$ $=\left(s_{t}^{2}\bm{U}_{k}^{\star}\bm{\Lambda}_{k}^{\star}\bm{U}_{k}^{\star T}+\gamma_{t}^{2}\bm{I}_{n}\right)^{-1}=\frac{1}{\gamma_{t}^{2}}\left(\bm{I}_{n}-\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\right),$
where $\bm{D}_{k}^{\star}=\mathrm{diag}\left(\frac{s_{t}^{2}\lambda_{k,1}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,1}^{\star}},\dots,\frac{s_{t}^{2}\lambda_{k,d_{k}}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,d_{k}}^{\star}}\right)$. This, together with the above equation, implies (26). ∎
Using the above result and (4), we compute $\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]$ when the data $\bm{x}_{0}$ is drawn from the MoLRG distribution as follows:
Lemma 1.
Suppose $\bm{x}_{0}$ is drawn from the MoLRG distribution with parameters $\{\pi_{k}\}_{k=1}^{K}$, $\{\bm{\mu}_{k}^{\star}\}_{k=1}^{K}$, and $\{\bm{\Sigma}^{\star}_{k}\}_{k=1}^{K}$. For each time $t\in(0,1]$, it holds that
$\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]=\sum_{k=1}^{K}w_{k}^{\star}(\bm{x}_{t})\left(\bm{\mu}_{k}^{\star}+\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\left(\frac{\bm{x}_{t}}{s_{t}}-\bm{\mu}_{k}^{\star}\right)\right),$
where
$\bm{D}_{k}^{\star}=\mathrm{diag}\left(\frac{s_{t}^{2}\lambda_{k,1}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,1}^{\star}},\dots,\frac{s_{t}^{2}\lambda_{k,d_{k}}^{\star}}{\gamma_{t}^{2}+s_{t}^{2}\lambda_{k,d_{k}}^{\star}}\right),\ w_{k}^{\star}(\bm{x}):=\frac{\pi_{k}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{k}^{\star},s_{t}^{2}\bm{\Sigma}_{k}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)}{\sum_{l=1}^{K}\pi_{l}\mathcal{N}\left(\bm{x};s_{t}\bm{\mu}_{l}^{\star},s_{t}^{2}\bm{\Sigma}_{l}^{\star}+\gamma_{t}^{2}\bm{I}_{n}\right)}.$
Proof.
According to (4) and (26) in Proposition 1, we compute
$\mathbb{E}\left[\bm{x}_{0}|\bm{x}_{t}\right]$ $=\frac{\bm{x}_{t}+\gamma_{t}^{2}\nabla\log p_{t}(\bm{x}_{t})}{s_{t}}=\sum_{k=1}^{k}w_{k}^{\star}(\bm{x}_{t})\left(\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\left(\frac{\bm{x}_{t}}{s_{t}}-\bm{\mu}_{k}^{\star}\right)+\bm{\mu}_{k}^{\star}\right).$
This lemma implies that the ground-truth posterior mean is a convex combination of the terms $\bm{\mu}_{k}^{\star}+\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\left({\bm{x}_{t}}/{s_{t}}-\bm{\mu}_{k}^{\star}\right)$, where the weights are $w_{k}^{\star}(\bm{x})$ for each $k\in[K]$.
Appendix B Proofs in Section 3
When $\bm{\mu}^{\star}_{k}=\bm{0}$ and $\bm{\Lambda}_{k}^{\star}=\bm{I}_{d_{k}}$ for each $k\in[K]$, we focus on a special instance of the MoLRG  distribution in Definition 1 as follows:
$\bm{x}_{0}\sim\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\bm{0},\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T}\right).$ (28)
This, together with Lemma 1, yields that the optimal parametrization for the DAE to learn the above distribution is
$\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)=\dfrac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\sum_{k=1}^{K}w_{k}(\bm{\theta};\bm{x}_{t})\bm{U}_{k}\bm{U}_{k}^{T}\bm{x}_{t},$ (29)
where $\bm{U}_{k}\in\mathcal{O}^{n\times d_{k}}$ for each $k\in[K]$ and $\phi_{t}=s_{t}^{2}/(2\gamma_{t}^{2}(s_{t}^{2}+\gamma_{t}^{2}))$. and
$w_{k}(\bm{\theta};\bm{x}_{t})=\frac{\pi_{k}\exp\left(\phi_{t}\|\bm{U}_{k}^{T}\bm{x}_{t}\|^{2}\right)}{\sum_{l=1}^{K}\pi_{l}(s_{t}^{2}+\gamma_{t}^{2})^{-\frac{d_{l}-d_{k}}{2}}\gamma_{t}^{d_{l}-d_{k}}\exp\left(\phi_{t}\|\bm{U}_{l}^{T}\bm{x}_{t}\|^{2}\right)}.$
B.1 Proof of Theorem 1
Proof of Theorem 1.
Plugging (11) into the integrand of (5) yields
$\ \mathbb{E}_{\bm{\epsilon}}\left[\left\|\frac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\bm{U}\bm{U}^{T}\left(s_{t}\bm{x}^{(i)}+\gamma_{t}\bm{\epsilon}\right)-\bm{x}^{(i)}\right\|^{2}\right]$
$=$ $\ \left\|\frac{s_{t}^{2}}{s_{t}^{2}+\gamma_{t}^{2}}\bm{U}\bm{U}^{T}\bm{x}^{(i)}-\bm{x}^{(i)}\right\|^{2}+\frac{(s_{t}\gamma_{t})^{2}}{(s_{t}^{2}+\gamma_{t})^{2}}\mathbb{E}_{\bm{\epsilon}}\left[\|\bm{U}\bm{U}^{T}\bm{\epsilon}\|^{2}\right]$
$=$ $\left\|\frac{s_{t}^{2}}{s_{t}^{2}+\gamma_{t}^{2}}\bm{U}\bm{U}^{T}\bm{x}^{(i)}-\bm{x}^{(i)}\right\|^{2}+\frac{(s_{t}\gamma_{t})^{2}d}{(s_{t}^{2}+\gamma_{t})^{2}},$
where the first equality follows from $\mathbb{E}_{\bm{\epsilon}}[\langle\bm{x},\bm{\epsilon}\rangle]=0$ for any given $\bm{x}\in\mathbb{R}^{n}$ due to $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, and the second equality uses $\mathbb{E}_{\bm{\epsilon}}\left[\|\bm{U}\bm{U}^{T}\bm{\epsilon}\|^{2}\right]=\mathbb{E}_{\bm{\epsilon}}\left[\|\bm{U}^{T}\bm{\epsilon}\|^{2}\right]=\sum_{i=1}^{d}\mathbb{E}_{\bm{\epsilon}}\left[\|\bm{u}_{i}^{T}\bm{\epsilon}\|^{2}\right]=d$ due to $\bm{U}\in\mathcal{O}^{n\times d}$ and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$. This, together with $\gamma_{t}=s_{t}\sigma_{t}$ and (5), yields
$\ell(\bm{U})=\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{1}\lambda_{t}\left(\|\bm{x}^{(i)}\|^{2}-\frac{1+2\sigma_{t}^{2}}{(1+\sigma_{t}^{2})^{2}}\|\bm{U}^{T}\bm{x}^{(i)}\|^{2}+\frac{\sigma_{t}^{2}d}{(1+\sigma_{t}^{2})^{2}}\right){\rm d}t,$
Obviously, minimizing the above function in terms of $\bm{U}$ amounts to
$\min_{\bm{U}^{T}\bm{U}=\bm{I}_{d}}-\int_{0}^{1}\frac{(1+2\sigma_{t}^{2})\lambda_{t}}{(1+\sigma_{t}^{2})^{2}}{\rm d}t\frac{1}{N}\sum_{i=1}^{N}\|\bm{U}^{T}\bm{x}^{(i)}\|^{2},$
which is equivalent to Problem (12). ∎
B.2 Proof of Theorem 2
Proof of Theorem 2.
For ease of exposition, let
$\bm{X}=\begin{bmatrix}\bm{x}^{(1)}&\dots&\bm{x}^{(N)}\end{bmatrix}\in\mathbb{R}^{n\times N},\ \bm{A}=\begin{bmatrix}\bm{a}_{1}&\dots&\bm{a}_{N}\end{bmatrix}\in\mathbb{R}^{d\times N},\ \bm{E}=\begin{bmatrix}\bm{e}_{1}&\dots&\bm{e}_{N}\end{bmatrix}\in\mathbb{R}^{n\times N}.$
Using this and (10), we obtain
$\bm{X}=\bm{U}^{\star}\bm{A}+\bm{E}.$ (30)
Let $r_{A}:=\mathrm{rank}(\bm{A})\leq\min\{d,N\}$ and $\bm{A}=\bm{U}_{A}\bm{\Sigma}_{A}\bm{V}_{A}^{T}$ be an singular value decomposition (SVD) of $\bm{A}$, where $\bm{U}_{A}\in\mathcal{O}^{d\times r_{A}}$, $\bm{V}_{A}\in\mathcal{O}^{N\times r_{A}}$, and $\bm{\Sigma}_{A}\in\mathbb{R}^{r_{A}\times r_{A}}$. It follows from Theorem 1 that Problem (5) with the parameterization (11) is equivalent to Problem (12).
(i) Suppose that $N\geq d$. Applying Lemma 2 with $\varepsilon={1}/{(2c_{1})}$ to $\bm{A}\in\mathbb{R}^{d\times N}$, it holds with probability at least $1-1/2^{N-d+1}-\exp\left(-c_{2}N\right)$ that
$\sigma_{\min}(\bm{A})=\sigma_{d}(\bm{A})\geq\frac{\sqrt{N}-\sqrt{d-1}}{2c_{1}},$ (31)
where $c_{1},c_{2}>0$ are constants depending polynomially only on the Gaussian moment. This implies $r_{A}=d$ and $\bm{U}_{A}\in\mathcal{O}^{d}$. Since Problem (12) is a PCA problem, the columns of any optimal solution $\hat{\bm{U}}\in\mathcal{O}^{n\times d}$ consist of left singular vectors associated with the top $d$ singular values of $\bm{X}$. This, together with Wedin’s Theorem [47] and (30), yields
$\left\|\hat{\bm{U}}\hat{\bm{U}}^{T}-\bm{U}^{\star}\bm{U}^{\star T}\right\|_{F}=\left\|\hat{\bm{U}}\hat{\bm{U}}^{T}-(\bm{U}^{\star}\bm{U}_{A})(\bm{U}^{\star}\bm{U}_{A})^{T}\right\|_{F}\leq\frac{2\|\bm{E}\|_{F}}{\sigma_{\min}(\bm{A})}=\frac{4c_{1}\|\bm{E}\|_{F}}{\sqrt{N}-\sqrt{d-1}}.$
This, together with absorbing $4$ into $c_{1}$, yields (13).
(ii) Suppose that $N<d$. According to Lemma 2 with $\varepsilon={1}/{(2c_{1})}$, it holds with probability at least $1-1/2^{d-N+1}-\exp\left(-c_{2}d\right)$ that
$\sigma_{\min}(\bm{A})=\sigma_{N}(\bm{A})\geq\frac{\sqrt{d}-\sqrt{N-1}}{2c_{1}},$ (32)
where $c_{1},c_{2}>0$ are constants depending polynomially only on the Gaussian moment. This implies $r_{A}=N$ and $\bm{U}_{A}\in\mathcal{O}^{d\times N}$. This, together with the fact that $\bm{A}=\bm{U}_{A}\bm{\Sigma}_{A}\bm{V}_{A}^{T}$ is an SVD of $\bm{A}$, yields that $\bm{U}^{\star}\bm{A}=\left(\bm{U}^{\star}\bm{U}_{A}\right)\bm{\Sigma}_{A}\bm{V}_{A}^{T}$ is an SVD of $\bm{U}^{\star}\bm{A}$ with $\bm{U}^{\star}\bm{U}_{A}\in\mathcal{O}^{n\times N}$. Note that $\mathrm{rank}(\bm{X})\leq N$. Let $\bm{X}=\bm{U}_{X}\bm{\Sigma}_{X}\bm{V}_{X}^{T}$ be an SVD of $\bm{X}$, where $\bm{U}_{X}\in\mathcal{O}^{n\times N}$, $\bm{V}_{X}\in\mathcal{O}^{N}$, and $\bm{\Sigma}_{X}\in\mathbb{R}^{N\times N}$. This, together with Wedin’s Theorem [47] and (32), yields
$\left\|\bm{U}_{X}\bm{U}_{X}^{T}-\bm{U}^{\star}\bm{U}_{A}\bm{U}_{A}^{T}\bm{U}^{\star T}\right\|_{F}\leq\frac{2\|\bm{E}\|_{F}}{\sigma_{\min}(\bm{A})}=\frac{4c_{1}\|\bm{E}\|_{F}}{\sqrt{d}-\sqrt{N-1}}.$ (33)
Note that Problem (12) has infinite optimal solutions when $N<d$, which take the form of
$\hat{\bm{U}}=\begin{bmatrix}\bm{U}_{X}&\bar{\bm{U}}_{X}\end{bmatrix}\in\mathcal{O}^{n\times d}.$
Now, we consider that $\bar{\bm{U}}_{X}\in\mathcal{O}^{n\times(d-N)}$ is an optimal solution of the following problem:
$\min_{\bm{V}\in\mathcal{O}^{n\times(d-N)},\bm{U}_{X}^{T}\bm{V}=\bm{0}}\|\bm{V}^{T}\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\|_{F}^{2}.$ (34)
Then, one can verify that the rank of the following matrix is at most $d$:
$\bm{B}:=\begin{bmatrix}\bm{U}_{X}&\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\end{bmatrix}$
Then, if $n\geq 2d-N$, it is easy to see that the optimal value of Problem (34) is 00. If $n<2d-N$, the optima value is achieved at $\bm{V}^{\star}=[\bm{V}_{1}^{\star}\ \bm{V}_{2}^{\star}]$ with $\bm{V}_{1}^{\star}\in\mathbb{R}^{n\times(n-d)}$ and $\bm{V}_{2}^{\star}\in\mathbb{R}^{n\times(2d-N-n)}$ satisfying $\bm{V}_{1}^{\star T}\bm{B}=\bm{0}$, which implies
$\|\bm{V}^{\star T}\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\|_{F}^{2}=\|\bm{V}_{2}^{\star T}\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\|_{F}^{2}\leq 2d-N-n.$
Consequently, the optimal value of Problem (34) is less than
$\max\left\{0,2d-(n+N)\right\}$ (35)
Then, we obtain that
$\left\|\hat{\bm{U}}\hat{\bm{U}}^{T}-\bm{U}^{\star}\bm{U}^{\star T}\right\|_{F}$ $=\left\|\bm{U}_{X}\bm{U}_{X}^{T}+\bar{\bm{U}}_{X}\bar{\bm{U}}_{X}^{T}-\bm{U}^{\star}\bm{U}_{A}\bm{U}_{A}^{T}\bm{U}^{\star T}-\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\bm{U}^{\star T}\right\|$
$\geq\|\bar{\bm{U}}_{X}\bar{\bm{U}}_{X}^{T}-\bm{U}^{\star}(\bm{I}-\bm{U}_{A}\bm{U}_{A}^{T})\bm{U}^{\star T}\|_{F}-\left\|\bm{U}_{X}\bm{U}_{X}^{T}-\bm{U}^{\star}\bm{U}_{A}\bm{U}_{A}^{T}\bm{U}^{\star T}\right\|_{F}$
$\geq\sqrt{2(d-N)-2\max\left\{0,2d-(n+N)\right\}}-\frac{4c_{1}\|\bm{E}\|_{F}}{\sqrt{d}-\sqrt{N-1}}$
$\geq\sqrt{2\min\{d-N,n-d\}}-\frac{4c_{1}\|\bm{E}\|_{F}}{\sqrt{d}-\sqrt{N-1}},$
where the second inequality follows from $\bar{\bm{U}}_{X}=\bm{V}^{\star}$ and (35). Then, we complete the proof.
B.3 Theoretical Justification of the DAE in (15)
Since $\bm{x}_{t}=s_{t}\bm{x}_{0}+\gamma_{t}\bm{\epsilon}$, we compute
$\mathbb{E}_{\bm{\epsilon}}\left[\|\bm{U}_{k}^{T}(s_{t}\bm{x}_{0}+\gamma_{t}\bm{\epsilon})\|^{2}\right]=s_{t}^{2}\|\bm{U}_{k}^{T}\bm{x}_{0}\|^{2}+\gamma_{t}^{2}\mathbb{E}_{\bm{\epsilon}}[\|\bm{U}_{k}^{T}\bm{\epsilon}\|^{2}]=s_{t}^{2}\|\bm{U}_{k}^{T}\bm{x}_{0}\|^{2}+\gamma_{t}^{2}d,$
where the first equality is due to $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ and $\mathbb{E}_{\bm{\epsilon}}[\langle\bm{U}_{k}^{T}\bm{x}_{0},\bm{U}_{k}^{T}\bm{\epsilon}\rangle]=\bm{0}$ for each $k\in[K]$. This implies that when $n$ is sufficiently large, we can approximate $w_{k}(\bm{\theta};\bm{x}_{t})$ in (29) well by
$w_{k}(\bm{\theta};\bm{x}_{t})$ $\approx\frac{\exp\left(\phi_{t}\left(s_{t}^{2}\|\bm{U}_{k}^{T}\bm{x}_{0}\|^{2}+\gamma_{t}^{2}d\right)\right)}{\sum_{l=1}^{K}\exp\left(\phi_{t}\left(s_{t}^{2}\|\bm{U}_{l}^{T}\bm{x}_{0}\|^{2}+\gamma_{t}^{2}d\right)\right)}.$
This soft-max function can be further approximated by the hard-max function. Therefore, we directly obtain (16).
B.4 Proof of Theorem 3
Equipped with the above setup, we are ready to prove Theorem 3.
Proof of Theorem 3.
Plugging (15) into the integrand of (5) yields
$\ \mathbb{E}_{\bm{\epsilon}}\left[\left\|\frac{s_{t}}{s_{t}^{2}+\gamma_{t}^{2}}\sum_{k=1}^{K}\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\bm{U}_{k}\bm{U}_{k}^{T}(s_{t}\bm{x}^{(i)}+\gamma_{t}\bm{\epsilon})-\bm{x}^{(i)}\right\|^{2}\right]$
$=$ $\ \left\|\frac{s_{t}^{2}}{s_{t}^{2}+\gamma_{t}^{2}}\sum_{k=1}^{K}\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\bm{U}_{k}\bm{U}_{k}^{T}\bm{x}^{(i)}-\bm{x}^{(i)}\right\|^{2}+\frac{(s_{t}\gamma_{t})^{2}}{(s_{t}^{2}+\gamma_{t}^{2})^{2}}\mathbb{E}_{\bm{\epsilon}}\left[\left\|\sum_{k=1}^{K}\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\bm{U}_{k}\bm{U}_{k}^{T}\bm{\epsilon}\right\|^{2}\right]$
$=$ $\ \frac{s_{t}^{2}}{s_{t}^{2}+\gamma_{t}^{2}}\sum_{k=1}^{K}\left(\frac{s_{t}^{2}}{s_{t}^{2}+\gamma_{t}^{2}}\hat{w}_{k}^{2}(\bm{\theta};\bm{x}^{(i)})-2\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\right)\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}+\|\bm{x}^{(i)}\|^{2}+\frac{(s_{t}\gamma_{t})^{2}d}{(s_{t}^{2}+\gamma_{t}^{2})^{2}}\sum_{k=1}^{K}\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)}),$
where the first equality follows from $\mathbb{E}_{\bm{\epsilon}}[\langle\bm{x},\bm{\epsilon}\rangle]=0$ for any fixed $\bm{x}\in\mathbb{R}^{n}$ due to $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, and the last equality uses $\bm{U}_{k}\in\mathcal{O}^{n\times d}$ and $\bm{U}_{k}^{T}\bm{U}_{l}=\bm{0}$ for all $k\neq l$. This, together with (5) and $\gamma_{t}=s_{t}\sigma_{t}$, yields
$\ell(\bm{\theta})$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\int_{0}^{1}\frac{\lambda_{t}}{1+\sigma_{t}^{2}}\left(\frac{1}{1+\sigma_{t}^{2}}\hat{w}_{k}^{2}(\bm{\theta};\bm{x}^{(i)})-2\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\right){\rm d}t\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}+$
$\frac{1}{N}\int_{0}^{1}\lambda_{t}{\rm d}t\sum_{i=1}^{N}\|\bm{x}^{(i)}\|^{2}+\left(\int_{0}^{1}\frac{\sigma_{t}^{2}\lambda_{t}}{(1+\sigma_{t}^{2})^{2}}{\rm d}t\right)\frac{d}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\hat{w}_{k}^{2}(\bm{\theta};\bm{x}^{(i)}).$
According to (15), we can partition $[N]$ into $\{C_{k}(\bm{\theta})\}_{k=1}^{K}$, where $C_{k}(\bm{\theta})$ for each $k\in[K]$ is defined as follows:
$C_{k}(\bm{\theta}):=\left\{i\in[N]:\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|\geq\|\bm{U}_{l}^{T}\bm{x}^{(i)}\|,\ \forall l\neq k\right\},\forall k\in[K].$ (36)
Then, we obtain
$\sum_{i=1}^{N}\sum_{k=1}^{K}\hat{w}_{k}^{2}(\bm{\theta};\bm{x}^{(i)})=\sum_{k=1}^{K}\sum_{i\in C_{k}(\bm{\theta})}1=N.$
This, together with plugging (36) into the above loss function, yields minimizing $\ell(\bm{\theta})$ is equivalent to minimizing
$\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\int_{0}^{1}\frac{\lambda_{t}}{1+\sigma_{t}^{2}}\left(\frac{1}{1+\sigma_{t}^{2}}\hat{w}_{k}^{2}(\bm{\theta};\bm{x}^{(i)})-2\hat{w}_{k}(\bm{\theta};\bm{x}^{(i)})\right){\rm d}t\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}$
$=$ $\left(\int_{0}^{1}\frac{\lambda_{t}}{1+\sigma_{t}^{2}}\left(\frac{1}{1+\sigma_{t}^{2}}-2\right){\rm d}t\right)\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}(\bm{\theta})}\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}.$
Since $\frac{\lambda_{t}}{1+\sigma_{t}^{2}}\left(\frac{1}{1+\sigma_{t}^{2}}-2\right)<0$ for all $t\in[0,1]$, minimizing the above function is equivalent to
$\max_{\bm{\theta}}\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in C_{k}(\bm{\theta})}\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2}\qquad\mathrm{s.t.}\ \left[\bm{U}_{1}\ \dots\ \bm{U}_{K}\right]\in\mathcal{O}^{n\times dK}.$
Then, we complete the proof. ∎
B.5 Proof of Theorem 4
Proof of Theorem 4.
For ease of exposition, let $\delta:=\max\{\|\bm{e}_{i}\|:i\in[N]\}$,
$f(\bm{\theta}):=\sum_{k=1}^{K}\sum_{i\in C_{k}(\bm{\theta})}\|\bm{U}_{k}^{T}\bm{x}^{(i)}\|^{2},$
and for each $k\in[K]$,
$C_{k}^{\star}:=\left\{i\in[N]:\bm{x}^{(i)}=\bm{U}_{k}^{\star}\bm{a}_{i}+\bm{e}_{i}\right\}.$
Suppose that (53) and (54) hold with $\bm{V}=\hat{\bm{U}}_{k}$ for all $i\in[N]$ and $k\neq l\in[K]$, which happens with probability $1-2K^{2}N^{-1}$ according to Lemma 4. This implies that for all $i\in[N]$ and $k\neq l\in[K]$,
$\sqrt{d}-(2\sqrt{\log N}+2)\leq\|\bm{a}_{i}\|\leq\sqrt{d}+(2\sqrt{\log N}+2),$ (37)
$\|\hat{\bm{U}}_{k}^{T}\bm{U}_{l}^{\star}\|_{F}-(2\sqrt{\log N}+2)\leq\|\hat{\bm{U}}_{k}^{T}\bm{U}_{l}^{\star}\bm{a}_{i}\|\leq\|\hat{\bm{U}}_{k}^{T}\bm{U}_{l}^{\star}\|_{F}+(2\sqrt{\log N}+2).$ (38)
Recall that the underlying basis matrices are denoted by $\bm{\theta}^{\star}=\{\bm{U}_{k}^{\star}\}_{k=1}^{K}$ and the optimal basis matrices are denoted by $\hat{\bm{\theta}}=\{\hat{\bm{U}}_{k}\}_{k=1}^{K}$.
First, we claim that $C_{k}(\bm{\theta}^{\star})=C_{k}^{\star}$ for each $k\in[K]$. Indeed, for each $i\in C_{k}^{\star}$, we compute
$\|\bm{U}_{k}^{\star T}\bm{x}^{(i)}\|=\|\bm{U}_{k}^{\star T}(\bm{U}_{k}^{\star}\bm{a}_{i}+\bm{e}_{i})\|=\|\bm{a}_{i}+\bm{U}_{k}^{\star T}\bm{e}_{i}\|\geq\|\bm{a}_{i}\|-\|\bm{e}_{i}\|,$ (39)
$\|\bm{U}_{l}^{\star T}\bm{x}^{(i)}\|=\|\bm{U}_{l}^{\star^{T}}(\bm{U}_{k}^{\star}\bm{a}_{i}+\bm{e}_{i})\|=\|\bm{U}_{l}^{\star T}\bm{e}_{i}\|\leq\|\bm{e}_{i}\|,\ \forall l\neq k.$ (40)
This, together with (37), $\|\bm{e}_{i}\|<(\sqrt{d}-2\sqrt{\log N})/2$, and $d\gtrsim\log N$, implies $\|\bm{U}_{k}^{\star T}\bm{x}_{i}\|\geq\|\bm{U}_{l}^{\star T}\bm{x}_{i}\|$ for all $l\neq k$. Therefore, we have $i\in C_{k}(\bm{\theta}^{\star})$ due to (36). Therefore, we have $C_{k}^{\star}\subseteq C_{k}(\bm{\theta}^{\star})$ for each $k\in[K]$. This, together with the fact that they respectively denote a partition of $[N]$, yields $C_{k}(\bm{\theta}^{\star})=C_{k}^{\star}$ for each $k\in[K]$. Now, we compute
$f(\bm{\theta}^{\star})$ $=\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\|\bm{U}_{k}^{\star T}\bm{x}^{(i)}\|^{2}=\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\|\bm{a}_{i}+\bm{U}_{k}^{\star T}\bm{e}_{i}\|^{2}$
$=\sum_{i=1}^{N}\|\bm{a}_{i}\|^{2}+2\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\langle\bm{a}_{i},\bm{U}_{k}^{\star T}\bm{e}_{i}\rangle+\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\|\bm{U}_{k}^{\star T}\bm{e}_{i}\|^{2}.$ (41)
Next, we compute
$f(\hat{\bm{\theta}})$ $=\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})}\|\hat{\bm{U}}_{k}^{T}\bm{x}^{(i)}\|^{2}=\sum_{l=1}^{K}\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})\cap C_{l}^{\star}}\|\hat{\bm{U}}_{k}^{T}(\bm{U}_{l}^{\star}\bm{a}_{i}+\bm{e}_{i}))\|^{2}$
$=\sum_{l=1}^{K}\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})\cap C_{l}^{\star}}\left(\|\hat{\bm{U}}_{k}^{T}\bm{U}_{l}^{\star}\bm{a}_{i}\|^{2}+2\langle\bm{a}_{i},\bm{U}_{l}^{\star T}\hat{\bm{U}}_{k}\hat{\bm{U}}_{k}^{T}\bm{e}_{i}\rangle\right)+\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})}\|\hat{\bm{U}}_{k}^{T}\bm{e}_{i}\|^{2}.$
This, together with $f(\hat{\bm{\theta}})\geq f(\bm{\theta}^{\star})$ and (B.5), yields
$\ \sum_{i=1}^{N}\|\bm{a}_{i}\|^{2}-\sum_{l=1}^{K}\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})\cap C_{l}^{\star}}\|\hat{\bm{U}}_{k}^{T}\bm{U}_{l}^{\star}\bm{a}_{i}\|^{2}\leq\sum_{l=1}^{K}\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})\cap C_{l}^{\star}}2\langle\bm{a}_{i},\bm{U}_{l}^{\star T}\hat{\bm{U}}_{k}\hat{\bm{U}}_{k}^{T}\bm{e}_{i}\rangle$
$\ +\sum_{k=1}^{K}\sum_{i\in C_{k}(\hat{\bm{\theta}})}\|\hat{\bm{U}}_{k}^{T}\bm{e}_{i}\|^{2}-2\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\langle\bm{a}_{i},\bm{U}_{k}^{\star T}\bm{e}_{i}\rangle-\sum_{k=1}^{K}\sum_{i\in C_{k}^{\star}}\|\bm{U}_{k}^{\star T}\bm{e}_{i}\|^{2}$
$\ \leq 4\delta\sum_{i=1}^{N}\|\bm{a}_{i}\|+N\delta^{2}\leq 6\delta N\sqrt{d}+N\delta^{2},$ (42)
where the second inequality follows from $\|\bm{e}_{i}\|\leq\delta$ for all $i\in[N]$ and $\bm{U}_{k}^{\star},\hat{\bm{U}}_{k}\in\mathcal{O}^{n\times d}$ for all $k\in[K]$, and the last inequality uses (37) and $\sqrt{d}\geq 4(\sqrt{\log N}+1)$ due to $d\gtrsim\log N$.
For ease of exposition, let $N_{kl}:=|C_{k}(\hat{\bm{\theta}})\cap C_{l}^{\star}|$. According to the pigeonhole principle, there exists a permutation $\pi:[K]\to[K]$ such that there exists $k\in[K]$ such that $N_{\pi(k)k}\geq N/K^{2}$. This, together with (B.5), yields
$6\delta N\sqrt{d}+N\delta^{2}$ $\geq\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\left(\|\bm{a}_{i}\|^{2}-\|\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star}\bm{a}_{i}\|^{2}\right)$
$=\left\langle\bm{I}-\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star},\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\bm{a}_{i}\bm{a}_{i}^{T}\right\rangle.$ (43)
According to Lemma 5 and $N_{\pi(k)k}\geq N/K^{2}$, it holds with probability at least $1-2K^{4}N^{-2}$ that
$\left\|\frac{1}{N_{\pi(k)k}}\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\bm{a}_{i}\bm{a}_{i}^{T}-\bm{I}\right\|\leq\frac{9(\sqrt{d}+\sqrt{\log(N_{\pi(k)k}})}{\sqrt{N_{\pi(k)k}}}.$
This, together with the Weyl’s inequality, yields
$\lambda_{\min}\left(\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\bm{a}_{i}\bm{a}_{i}^{T}\right)$ $\geq N_{\pi(k)k}-9\sqrt{N_{\pi(k)k}}\left(\sqrt{d}+\sqrt{\log(N_{\pi(k)k})}\right)$
$\geq\frac{N}{K^{2}}-\frac{9\sqrt{N}}{K}\left(\sqrt{d}+\sqrt{\log N}\right)\geq\frac{N}{2K^{2}},$
where the second inequality follows from $N/K^{2}\leq N_{\pi(k)k}\leq N$ and the last inequality is due to $\sqrt{N}\geq 18K(\sqrt{d}+\sqrt{\log N})$. Using this and Lemma 6, we obtain
$\left\langle\bm{I}-\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star},\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\bm{a}_{i}\bm{a}_{i}^{T}\right\rangle$ $\geq\lambda_{\min}\left(\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{k}^{\star}}\bm{a}_{i}\bm{a}_{i}^{T}\right)\mathrm{Tr}\left(\bm{I}-\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star}\right)$
$\geq\frac{N}{2K^{2}}\mathrm{Tr}\left(\bm{I}-\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star}\right).$
This, together with (B.5), implies
$\mathrm{Tr}\left(\bm{I}-\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star}\right)\leq 2K^{2}\left(6\delta\sqrt{d}+\delta^{2}\right).$
Using this and $[\bm{U}_{1}^{\star},\dots,\bm{U}_{k}^{\star}]\in\mathcal{O}^{n\times dK}$, we obtain
$\sum_{l\neq k}\|\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{l}^{\star}\|_{F}^{2}$ $=\mathrm{Tr}\left(\sum_{l\neq k}\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{l}^{\star}\bm{U}_{l}^{\star T}\hat{\bm{U}}_{\pi(k)}\right)\leq\mathrm{Tr}\left(\bm{I}-\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T}\hat{\bm{U}}_{\pi(k)}\right)$
$\leq 2K^{2}\left(6\delta\sqrt{d}+\delta^{2}\right)\leq\frac{3d}{4},$ (44)
where the last inequality follows $\delta\leq\sqrt{d}/(24K^{2})$. According to (B.5), we have
$6\delta N\sqrt{d}+N\delta^{2}$ $\geq\sum_{l\neq k}^{K}\sum_{i\in C_{\pi(k)}(\hat{\bm{\theta}})\cap C_{l}^{\star}}\left(\|\bm{a}_{i}\|^{2}-\|\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{l}^{\star}\bm{a}_{i}\|^{2}\right)$
$\geq\sum_{l\neq k}^{K}N_{\pi(k)l}\left((\sqrt{d}-\alpha)^{2}-\left(\|\hat{\bm{U}}_{\pi(k)}^{T}\bm{U}_{l}^{\star}\|_{F}+\alpha\right)^{2}\right)\geq\frac{d}{8}\sum_{l\neq k}^{K}N_{\pi(k)l},$
where the second inequality uses (37) and (38), and the last inequality follows from (B.5) and $d\gtrsim\log N$. Therefore, we have for each $k\in[K]$,
$\sum_{l\neq k}^{K}N_{\pi(k)l}\leq\frac{48\delta N\sqrt{d}+8\delta^{2}N}{d}<1,$
where the last inequality uses $\delta\lesssim\sqrt{d/N}$. This implies $N_{\pi(l)k}=0$ for all $l\neq k$, and thus $C_{\pi(k)}(\hat{\bm{\theta}})\subseteq C_{k}^{\star}$. Using the same argument, we can show that $C_{\pi(l)}(\hat{\bm{\theta}})\subseteq C_{l}^{\star}$ for each $l\neq k$. Therefore, we have $C_{\pi(k)}(\hat{\bm{\theta}})=C_{k}^{\star}$ for each $k\in[K]$. In particular, using the union bound yields event holds with probability at least $1-2K^{2}N^{-1}$. Based on the above optimal assignment, we further show:
(i) Suppose that $N_{k}\geq d$ for each $k\in[K]$. This, together with (i) in Theorem 2 and $N_{k}\geq d$, yields (18).
(ii) Suppose that there exists $k\in[K]$ such that $N_{k}<d$. This, together with (ii) in Theorem 2 and $N_{k}\geq d$, yields (19).
Finally, applying the union bound yields the probability of these events. ∎
Appendix C Experimental Setups in Section 2
[…]
We set $K=10$, $d_{k}=20$ for MNIST and FashionMNIST, $K=10$, $d_{k}=200$ for CIFAR-10, and $K=5$, $d_{k}=500$ for FFHQ. Since FFHQ lacks annotated labels, we apply the expectation-maximization algorithm for clustering and label generation. For comparison, we use a Gaussian distribution with its mean and covariance set to the mean and covariance of all training samples. In addition, we train an EDM-based diffusion model [35] on each dataset as a comparison. We employ the second-order Heun Solver [35] with 35 steps as the diffusion sampler to numerically solve (3) and generate samples from learned distributions. For qualitative evaluation, we visualize samples from 12 initial noise inputs per dataset for both theoretical (MoLRG and Gaussian) and real diffusion models. For quantitative evaluation, we generate 10K noise samples and compute the Euclidean distance between the theoretical and real model outputs (defined in (8)).
Appendix D Experimental Setups in Section 3
In this section, we provide detailed setups for the experiment in Section 3.3. This experiment aims to validate the Theorem 2 and Theorem 4. Here, we present the stochastic gradient descent (SGD) algorithm for solving Problem (5) in Algorithm 1.
Algorithm 1 SGD for optimizing the training loss (5)
Input: Training samples $\{\bm{x}^{(i)}\}_{i=1}^{N}$
for $j=0,1,2,\dots,J$ do
1. Randomly select $\{(i_{m},t_{m})\}_{m=1}^{M}$, where $i_{m}\in[N]$ and $t_{m}\in(0,1)$ and a noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
2. Take a gradient step
$\bm{\theta}^{j+1}\leftarrow\bm{\theta}^{j}-\frac{\eta}{M}\sum_{m\in[M]}\nabla_{\bm{\theta}}\left\|\bm{x}_{\bm{\theta}^{j}}(s_{t_{m}}\bm{x}^{(i_{m})}+\gamma_{t_{m}}\bm{\epsilon},t_{m})-\bm{x}^{(i_{m})}\right\|^{2}$
end for
Now, we specify how to choose the parameters of the SGD in our implementation. We divide the time interval $[0,1]$ into $64$ time steps. When $K=1$, we set the learning rate $\eta=10^{-4}$, batch size $M=128N_{k}$, and number of iterations $J=10^{4}$. When $K=2$, we set the learning rate $\eta=2\times 10^{-5}$, batch size $M=1024$, number of iterations $J=10^{5}$. In particular, when $K=2$, we use the following tailor-designed initialization $\bm{\theta}^{0}=\{\bm{U}_{k}^{0}\}$ to improve the convergence of the SGD:
$\bm{U}_{k}^{0}=\bm{U}_{k}^{\star}+0.2\bm{\Delta},\ k\in\{1,2\},$ (45)
where $\bm{\Delta}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$. We calculate the success rate as follows. If the returned subspace basis matrices $\{\bm{U}_{k}\}_{k=1}^{K}$ satisfy
$\frac{1}{K}\mathbb{\sum}_{k=1}^{K}||\bm{U}_{\Pi(k)}\bm{U}_{\Pi(k)}^{T}-\bm{U}_{k}^{\star}\bm{U}_{k}^{\star T}||\leq 0.5$
for some permutation $\Pi:[K]\to[K]$, it is considered successful.
Appendix E Experimental Setups in Section 4
In this section, we provide detailed setups for the experiments in Section 4. Specifically, we describe the settings for using a U-Net-based diffusion model to (1) learn MoLRG distribution (Appendix E.1), (2) learn real-world image distribution (Appendix E.2), and (3) estimate the intrinsic dimension of real-world image distribution (Appendix E.3).
E.1 Learning the MoLRG distribution with U-Net
In our implementation, we set $\mathrm{ID}\in\{8,10,12\}$. When $\mathrm{ID}=8$, $N_{k}\in\{20,50,70,100,200,300,1000\}$; when $\mathrm{ID}=10$, $N_{k}\in\{100,150,200,250,300,1000\}$; when $\mathrm{ID}=12$, $N_{k}\in\{100,150,200,250,300,350,400,1000\}$. To train U-Net, we use the stochastic gradient descent in Algorithm 1. We use DDPM++ architecture [6] for the U-Net and EDM [35] noise scheduler. We set the learning rate $10^{-3}$, batch size $64$, and number of iterations $J=10^{4}$.
For a specific MoLRG distribution $p_{\text{data}}$ with $N$ pre-selected training data $\bm{x}^{(i)}\sim p_{\text{data}}$, the threshold $\delta$ is chosen such that the following inequality holds:
$\frac{1}{M_{z}}\sum_{j=1}^{M_{z}}\mathbb{I}\left(\min_{i\in[N]}\|\bm{\Psi}\left(\bm{x}^{(i)}\right)-\bm{\Psi}\left(\bm{z}^{(j)}\right)\|\geq\delta\right)=0.95.$ (46)
Intuitively, this definition ensures that with $95\%$ probability, a newly drawn sample $\bm{z}^{(j)}\sim p_{\text{data}}$ will be at least $\delta$ away (in the $\bm{\Psi}$-transformation space) from its nearest neighbor among the training samples $\bm{x}_{i}$. While (46) has a theoretical analytical solution, we approximate $\delta$ numerically in practice. Specifically, we set $M_{z}=10^{3}$, compute the minimum distance $\min_{i\in[N]}||\bm{\Psi}\left(\bm{x}\right)-\bm{\Psi}\left(\bm{y}_{i}\right)||_{2}$ for each, and set $\delta$ as the 5th percentile (i.e., the 0.05-quantile) of the resulting distance distribution. To empirically estimate (21), we set $M=10^{3}$.
To quantitatively estimate the transition in Figure 4 (top-left), we fit the curve using the following sigmoid-parameterized function:
$\mathrm{GL}\left(\frac{N}{\mathrm{ID}}\right)\approx f_{\texttt{MoLRG}}\left(\frac{N}{\mathrm{ID}}\right)=\frac{1}{1+\exp\left(-a(\log_{2}\left(N/\mathrm{ID}\right)-b)\right)},$ (47)
where the fitted parameters are $a=6.22$ and $b=5.20$. And we could numerically solved that $f_{\texttt{MoLRG}}^{-1}\left(0.95\right)=50.2$, indicating that U-Net architectures training on MoLRG distribution generalize when $N_{k}\geq 50.2d_{k}$. We use the same parameterized function (47) for the fitted curve in Figure 4 (bottom-left), by changing the input variable to $N/\mathrm{ID}^{2}$.
E.2 Learning real-world image data distributions with U-Net
To train diffusion models for real-world image datasets, we use the DDPM++ architecture [6] for U-Net and variance preserving (VP) [6] noise scheduler. The U-Net is trained using the Adam optimizer [59], a variant of SGD in Algorithm 1. We set the learning rate $\eta=10^{-3}$, batch size $M=512$, and the total number of iterations $10^{5}$. To empirically estimate (21), we set $M=10^{4}$.
The curve $f_{\texttt{real}}$ is parameterized the same way as $f_{\texttt{MoLRG}}$ in (47), with $a=1.88$ and $b=7.74$. Then, we numerically solve that $f^{-1}_{\texttt{real}}\left(0.95\right)=630.3$, indicating that U-Net architectures trained on real data distribution generalize when $N\geq 630.3\mathrm{ID}$. We use the same parameterized function (47) for the fitted curve in Figure 4 (bottom-right), by changing the input variable to $N/\mathrm{ID}^{2}$.
(a) PCA
(b) Diffusion model
Figure 6: Phase transition of learning the MoLRG distribution when $K=3$. The $x$-axis is the number of training samples and $y$-axis is the dimension of subspaces. We apply a subspace clustering method and train diffusion models for solving Problems (17) and (5), visualizing the results in (a) and (b), respectively.
E.3 Estimating the intrinsic dimension of real-world dataset
Figure 7: Low-rank property of the denoising autoencoder of trained diffusion models. We plot the numerical rank of the Jacobian of the denoising autoencoder, i.e., $\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$, against the timestep $t$ by training diffusion models on different datasets. We train diffusion models on image datasets CIFAR-10, CelebA, FFHQ, and AFHQ. The experimental details are provided in Appendix E.3.
In this subsection, we conduct numerical experiments to estimate the intrinsic dimension of real-world image data distribution. Following from Lemma 1, as $t\rightarrow 1$, we have
$\nabla_{\bm{x}_{t}}\mathbb{E}[\bm{x}_{t}|\bm{x}_{0}]\approx\frac{1}{s_{t}}\sum_{k=1}^{K}\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T},$ (48)
given $w_{k}^{*}(\bm{x}_{t})\approx 1$ and $\nabla_{\bm{x}_{t}}w_{k}^{*}(\bm{x}_{t})\approx\bm{0}$. This relationship allows us to estimate the intrinsic dimension $\mathrm{ID}$ of a MoLRG distribution as:
$\mathrm{ID}\coloneqq\mathrm{rank}\left(\sum_{k=1}^{K}\bm{U}_{k}^{\star}\bm{D}_{k}^{\star}\bm{U}_{k}^{\star T}\right)\approx\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\mathbb{E}[\bm{x}_{t}|\bm{x}_{0}]\right),$ (49)
when $t\rightarrow 1$. Note that the DAE $\bm{x}_{\bm{\theta}}(\cdot,t)$ of the trained diffusion models satisfies $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\approx\mathbb{E}[\bm{x}_{t}|\bm{x}_{0}]$. Combining this with the observation in Section 2.4 that real-world image distributions can be well approximated by MoLRG distributions, we conclude that the intrinsic dimension can be estimated by:
$\mathrm{ID}\approx\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right).$ (50)
We evaluate the intrinsic dimension $\mathrm{ID}$ over four different datasets: CIFAR-10, CelebA, FFHQ, and AFHQ. We resize images from FFHQ and AFHQ such that $n=3072$ for all datasets. we calculate the numerical rank of Jacobian $\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ through
$\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right):=\operatorname{\arg\min}\left\{r\in[1,n]:\frac{\sum_{i=1}^{r}\sigma_{i}^{2}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right)}{\sum_{i=1}^{n}\sigma_{i}^{2}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right)}>\eta^{2}\right\},$ (51)
with $\eta=0.99$ and recall $\sigma_{i}^{2}\left(\bm{A}\right)$ denotes the $i$-th singular value of matrix $\bm{A}$.
To select a timestep $t$ to estimate the intrinsic dimension, we evaluate $\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right)$ at different timesteps across multiple datasets, as shown in Figure 7. Specifically, Given a random initial noise $\bm{x}_{1}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, we use the diffusion model to generate a sequence of images $\{\bm{x}_{t}\}$ according to the reverse ODE in (3). Along the sampling trajectory $\{\bm{x}_{t}\}$, we estimate $\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right)$ at each timestep. For the experiments, we utilize the Elucidating Diffusion Model (EDM) with the EDM noise scheduler [35] and DDPM++ architecture [58]. Moreover, we employ an 18-step Heun’s solver for sampling and present the results for 12 of these steps ($t=0,0.001,0.003,0.027,0.171,0.277,0.494,0.650,0.712,0.815,0.858,0.934$). For each dataset, we random sample 15 initial noise $\bm{x}_{1}$, calculate the mean of $\mathrm{rank}(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t))$ along the trajectory $\{\bm{x}_{t}\}$.
As shown in Figure 7, the plot of rank against $t$ exhibits a U-shaped curve, with the lowest rank consistently occurring around $t=0.815$ across all datasets. Timesteps too close to $t=0$ or $t=1$ are unsuitable for estimating the intrinsic dimension: when $t$ approaches $1$, $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ becomes less accurate because the training loss (5) assigns a small weight $\lambda_{t}$ to such timesteps; when $t$ approaches 00, [35] parameterize $\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)$ to be $\bm{x}_{t}$, causing $\mathrm{rank}\left(\nabla_{\bm{x}_{t}}\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\right)=n$ to be naturally very high. Thus, we select $t=0.815$, the timestep that achieves the lowest rank, to estimate the intrinsic dimension of real-world datasets. The estimated $\mathrm{ID}$ is shown in Table 2.
Figure 8: Correspondence between the singular vectors of the Jacobian of the DAE and semantic image attributes. (a,c) Additional examples when $t=0.7$. (b) Ablation studies when $t=0.1$ and $0.9$.
Appendix F Auxiliary Results
First, we present a probabilistic result to prove Theorem 2, which provides an optimal estimate of the small singular values of a matrix with i.i.d. Gaussian entries. This lemma is proved in [60, Theorem 1.1] for subgaussian random variables. Note that a random variable $\xi$ is called subgaussian if there exists $c>0$ such that for all $t>0$,
$\mathbb{P}\left(|\xi|>t\right)\leq 2\exp\left(-\frac{t^{2}}{c^{2}}\right).$
We say that the minimal $c$ in this inequality is the subgaussian moment of $\xi$.
Lemma 2.
Let $\bm{A}$ be an $m\times n$ random matrix, where $m\geq n$, whose elements are independent copies of a subgaussian random variable with mean zero and unit variance. It holds for every $\varepsilon>0$ that
$\mathbb{P}\left(\sigma_{\min}(\bm{A})\geq\varepsilon(\sqrt{m}-\sqrt{n-1})\right)\geq 1-\left(c_{1}\varepsilon\right)^{m-n+1}-\exp\left(-c_{2}m\right),$
where $c_{1},c_{2}>0$ are constants depending polynomially only on the subgaussian moment.
Next, we present a probabilistic bound on the deviation of the norm of a weighted sum of squared Gaussian random variables from its mean. This is a direct extension of [61, Theorem 5.2.2].
Lemma 3.
Let $\bm{x}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ be a Gaussian random vector and $\lambda_{1},\dots,\lambda_{d}>0$ be constants. It holds for any $t>0$ that
$\mathbb{P}\left(\left|\sqrt{\sum_{i=1}^{d}\lambda_{i}^{2}x_{i}^{2}}-\sqrt{\sum_{i=1}^{d}\lambda_{i}^{2}}\right|\geq t+2\lambda_{\max}\right)\leq 2\exp\left(-\frac{t^{2}}{2\lambda_{\max}^{2}}\right),$ (52)
where $\lambda_{\max}=\max\{\lambda_{i}:i\in[d]\}$.
Based on the above lemma, we can further show the following concentration inequalities to estimate the norm of the standard norm Gaussian random vector.
Lemma 4.
Suppose that $\bm{a}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\bm{0},\bm{I}_{d})$ is a Gaussian random vector for each $i\in[N]$. The following statements hold:
(i) It holds for all $i\in[N]$ with probability at least $1-N^{-1}$ that
$\left|\|\bm{a}_{i}\|-\sqrt{d}\right|\leq 2\sqrt{\log N}+2.$ (53)
(ii) Let $\bm{V}\in\mathcal{O}^{n\times d}$ be given. For all $i\in C_{k}^{\star}$ and all $k\in[K]$, it holds with probability at least $1-2N^{-1}$ that
$\left|\|\bm{V}^{T}\bm{U}_{k}^{\star}\bm{a}_{i}\|-\|\bm{V}^{T}\bm{U}_{k}^{\star}\|_{F}\right|\leq 2\sqrt{\log N}+2.$ (54)
Proof.
(i) Applying Lemma 3 to $\bm{a}_{i}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$, together with setting $t=2\sqrt{\log N}$ and $\lambda_{j}=1$ for all $j\in[d]$, yields
$\mathbb{P}\left(\left|\|\bm{a}_{i}\|-\sqrt{d}\right|\geq 2\sqrt{\log N}+2\right)\leq 2N^{-2}.$
This, together with the union bound, yields that (53) holds with probability $1-N^{-1}$.
(ii) Let $\bm{V}^{T}\bm{U}^{\star}_{k}=\bm{P}\bm{\Sigma}\bm{Q}^{T}$ be a singular value decomposition of $\bm{V}^{T}\bm{U}^{\star}_{k}$, where $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ with the diagonal elements $0\leq\sigma_{d}\leq\dots\sigma_{1}\leq 1$ being the singular values of $\bm{V}^{T}\bm{U}^{\star}_{k}$ and $\bm{P},\bm{Q}\in\mathcal{O}^{d}$. This, together with the orthogonal invariance of the Gaussian distribution, yields
[…]
Using Lemma 3 with setting $t=2\sigma_{1}\sqrt{\log N}$ and $\lambda_{j}=\sigma_{j}\leq 1$ for all $j$ yields
$\mathbb{P}\left(\left|\|\bm{V}^{T}\bm{U}^{\star}_{k}\bm{a}_{i}\|-\|\bm{V}^{T}\bm{U}^{\star}_{k}\|_{F}\right|\geq\sigma_{1}\alpha\right)$ $=\mathbb{P}\left(\left|\sqrt{\sum_{j=1}^{d}\sigma_{j}^{2}a_{ij}^{2}}-\sqrt{\sum_{j=1}^{d}\sigma_{j}^{2}}\right|\geq\sigma_{1}\alpha\right)\leq 2N^{-2}.$
This, together with $\sigma_{1}\leq 1$ and the union bound, yields (54). ∎
Next, We present a spectral bound on the covariance estimation for the random vectors generated by the normal distribution.
Lemma 5.
Suppose that $\bm{a}_{1},\dots,\bm{a}_{N}\in\mathbb{R}^{d}$ are i.i.d. standard normal random vectors, i.e., $\bm{a}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\bm{0},\bm{I}_{d})$ for all $i\in[N]$. Then, it holds with probability at least $1-2N^{-2}$ that
$\left\|\frac{1}{N}\sum_{i=1}^{N}\bm{a}_{i}\bm{a}_{i}^{T}-\bm{I}_{d}\right\|\leq\frac{9(\sqrt{d}+\sqrt{\log N})}{\sqrt{N}},$ (56)
Proof.
According to [61, Theorem 4.7.1], it holds that
$\mathbb{P}\left(\left\|\frac{1}{N}\sum_{i=1}^{N}\bm{a}_{i}\bm{a}_{i}^{T}-\bm{I}_{d}\right\|\geq\frac{9(\sqrt{d}+\eta)}{\sqrt{N}}\right)\leq 2\exp\left(-2\eta^{2}\right),$
where $\eta>0$. Plugging $\eta=\sqrt{\log N}$ into the above inequality yields
$\mathbb{P}\left(\left\|\frac{1}{N}\sum_{i=1}^{N}\bm{a}_{i}\bm{a}_{i}^{T}-\bm{I}_{d}\right\|\geq\frac{9(\sqrt{d}+\sqrt{\log N})}{\sqrt{N}}\right)\leq 2N^{-2}.$
This directly implies (56). ∎
Lemma 6.
Let $\bm{A},\bm{B}\in\mathbb{R}^{n\times n}$ be positive semi-definite matrices. Then, it holds that
$\langle\bm{A},\bm{B}\rangle\geq\lambda_{\min}(\bm{A})\mathrm{Tr}(\bm{B}).$ (57)
Proof.
Let $\bm{U}\bm{\Lambda}\bm{U}^{T}=\bm{A}$ be an eigenvalue decompositon of $\bm{A}$, where $\bm{U}\in\mathcal{O}^{n}$ and $\bm{\Sigma}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ is a diagonal matrix with diagonal entries $\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0$ being the eigenvalues. Then, we compute
$\langle\bm{A},\bm{B}\rangle=\langle\bm{U}\bm{\Lambda}\bm{U}^{T},\bm{B}\rangle=\langle\bm{\Lambda},\bm{U}\bm{B}\bm{U}^{T}\rangle\geq\lambda_{\min}(\bm{A})\mathrm{Tr}(\bm{U}\bm{B}\bm{U}^{T})=\lambda_{\min}(\bm{A})\mathrm{Tr}(\bm{B}),$
where the inequality follows from $\lambda_{i}\geq 0$ for all $i\in[N]$ and $\bm{B}$ is a positive semidefinite matrix. ∎